### v1

# Look Twice before You Leap: A Rational Agent Framework for Localized Adversarial Anonymization

Abstract
Current LLM-based text anonymization frameworks usually rely on remote API services from powerful LLMs, which creates an inherent "privacy paradox": users must somehow disclose data to untrusted third parties for superior privacy preservation. Moreover, directly migrating these frameworks to local small-scale models (LSMs) offers a suboptimal solution with catastrophic collapse in utility based on our core findings. Our work argues that this failure stems not merely from the capability deficits of LSMs, but from the inherent irrationality of the greedy adversarial strategies employed by current state-of-the-art (SoTA) methods. We model the anonymization process as a trade-off between Marginal Privacy Gain (MPG) and Marginal Utility Cost (MUC), and demonstrate that greedy strategies inevitably drift into an irrational state. To address this, we propose Rational Localized Adversarial Anonymization (RLAA), a fully localized and training-free framework featuring an Attacker-Arbitrator-Anonymizer (A-A-A) architecture. RLAA introduces an arbitrator that acts as a rationality gatekeeper, validating the attacker’s inference to filter out feedback providing negligible benefits on privacy preservation. This mechanism enforces a rational early-stopping criterion, and systematically prevents utility collapse. Extensive experiments on different datasets demonstrate that RLAA achieves the best privacy utility trade-off, and in some cases even outperforms SoTA on the Pareto principle. Our code and datasets will be released upon acceptance.
Look Twice before You Leap: A Rational Agent Framework for
Localized Adversarial Anonymization
Donghang Duan    Xu Zheng University of Electronic Science and Technology of China
1 Introduction
With the proliferation of large language models (LLMs) adopted in various fields, processing textual data that potentially contain sensitive Personal Identifiable Information (PII) has become inevitable and ubiquitous deusser2025survey. This trend makes text anonymization crucial, not only to meet stringent regulatory requirements such as GDPR and CCPA albrecht2016gdpr; bonta2022california, but also to protect personal privacy while unlocking the value of data. However, this situation presents a tricky trade-off between semantic utility and privacy where over-anonymization destroys the semantic value while under-anonymization invites re-identification risks, which is particularly acute in sensitive domains like smart healthcare and legal technology morris2025diri; im2024exploring; liu2025legal. The rapid evolution of LLM has exacerbated this situation because of their dual roles as context-aware anonymizers and superior attackers staab2024large; wang2025unique. Consequently, traditional NER-based anonymization methods are rendered obsolete, shifting the paradigm towards complex semantic obfuscation that conceals latent identity with less sacrifice of utility loiseau2025tau.
To address this situation, advanced anonymization frameworks based on LLMs have emerged. As the current state-of-the-art (SoTA) method, Feedback-guided Adversarial Anonymization (FgAA) staab2024large employs a dynamic adversarial game: an attacker model attempts to re-identify personal information from the text generated by an anonymizer model, and the anonymizer uses the inference feedback to instruct the anonymization process for refinement. However, the superior performance of FgAA heavily relies on the capabilities of powerful or closed-source LLMs (e.g., GPT-4 achiam2023gpt or DeepSeek-V3.2-Exp deepseekai2024deepseekv32 in our research), which are typically accessible only via third-party APIs. This inherent dependency creates a fundamental "privacy paradox": to protect sensitive data, users must first send them to an external and uncontrolled commercial entity, which constitutes a critical privacy vulnerability smith2023identifying. This problem causes trust failure feretzakis2024trustworthy, and renders the architecture unacceptable in high-sensitivity scenarios.
Figure 1: Utility Failure of FgAA’s Naive Migration.
The most intuitive solution to this privacy paradox is to migrate adversarial frameworks to local small-scale models (LSMs). However, our empirical findings reveal that this naive migration results in a catastrophic utility collapse. As illustrated in Figure 1, the naive migration on Llama3-8B causes destructive over-editing. It indiscriminately strips away non-sensitive context and stylistic nuances, reducing the rich original narrative to a generic and vacuous summary ("The days are still memorable"). Besides, it is observed that this collapse persists even when implementing instruction refinement (e.g., outputting "unknown" for inferences with low confidence), which demonstrates that heuristic constraints alone cannot rectify the inherent irrationality of the greedy decision-making process.
We argue that this failure stems not merely from the capability deficits of LSMs, but from the inherent irrationality of the greedy adversarial strategies employed by current methods. Adopting an economic perspective, the anonymization process constitutes a trade-off between Marginal Privacy Gain (MPG) and Marginal Utility Cost (MUC). By quantifying this ratio as the Marginal Rate of Substitution ($\text{MRS}=\frac{\text{MUC}}{\text{MPG}}$), the root cause of the utility collapse becomes mathematically evident: Greedy strategies inevitably drift into an irrational state due to LSMs’ hallucinations and the law of diminishing returns, where the model executes destructive edits for non-existent or negligible leaks (see theoretical proofs in Appendix A). In this case, MUC continues to accumulate while MPG approaches zero, pushing the MRS to inefficient levels. A concurrent study called SEAL kim2025self aims to distill knowledge from a powerful LLM to an LSM by using Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) rafailov2023direct to avoid disclosing raw data to third-party APIs. However, it requires additional resources for post-training and high-quality synthetic datasets remain scarce and difficult to construct effectively yukhymenko2024synthetic. Besides, our experimental results in Section 4.2 shows that it still fails to handle the rationality of anonymization.
To address the irrationality of naive migration and the data scarcity of distillation methods, we propose RLAA (Rational Localized Adversarial Anonymization), a training-free framework designed for fully localized deployment. The core idea of RLAA is to compensate for the capability deficits of LSMs through architectural rationality rather than finetuning. RLAA features a novel "Attacker-Arbitrator-Anonymizer" (A-A-A) architecture, where the arbitrator functions as a rationality gatekeeper. Instead of blindly accepting feedback, this component compels the system to scrutinize the attacker’s suggestions before executing text modifications. Drawing on the idea of meta-reasoning alazraki2025meta; wen2025reinforcement, the arbitrator implements a dual-validation mechanism: it first filters hallucinations to ensure factual correctness, and simultaneously enforces a rational stopping criterion based on the marginal rate of substitution (MRS). Crucially, this mechanism prevents utility collapse by rejecting destructive edits driven by invalid feedback. Furthermore, it maintains robust privacy protection by focusing the modification budget solely on genuine leaks, effectively eliminating the deadweight loss of irrational editing without compromising valid defense. We conducted a comprehensive evaluation spanning baseline comparisons, component ablations, and generalization stress tests across mainstream LLMs (Llama3-8B llama3modelcard, Qwen2.5-7B qwen2.5) and DeepSeek-V3.2-Exp deepseekai2024deepseekv32. Additionally, we validated the agent’s rational decision-making process through a quantitative economic efficiency analysis based on the MRS metric. The results empirically demonstrate that RLAA structurally prevents utility collapse and achieves a superior privacy-utility trade-off compared to our baselines.
In summary, the main contributions of this study are as follows:
  • It is argued that migrating adversarial anonymization to local environments is crucial for eliminating API-based privacy risks, but our research shows that a naive migration to LSMs results in utility failure. This failure is analyzed and attributed to the inherent irrationality of greedy adversarial strategies via a marginal economic analysis in our work.
  • RLAA is proposed as a localized and training-free framework featuring an Attacker-Arbitrator-Anonymizer (A-A-A) architecture. By introducing an arbitrator inspired by meta-reasoning, it structurally enforces rational decision-making to prevent utility collapse.
  • Extensive experiments demonstrate that RLAA achieves a superior privacy-utility trade-off compared to our baselines. Notably, it achieves strict Pareto dominance on the reddit-self-disclosure dataset and demonstrate its mechanism generalization across different base models.
2 Related Work
2.1 Inference Attacks through LLMs
With increased reasoning capabilities, LLMs have evolved into potent privacy attackers capable of automatically inferring personal attributes from unstructured text staab24beyond; wang2025unique. staab24beyond demonstrated that models can accurately deduce fine-grained PII from casual online comments. Extending to high-stakes domains, nyffenegger2024anonymity revealed that LLMs could re-identify individuals in court decisions. Recent audits by panda2025daiq further confirmed that this inference capability is consistent across various model architectures. These studies show that LLM significantly lowers the threshold for privacy violations, enabling re-identification attacks to occur on an unprecedented scale staab24beyond.
2.2 Text Anonymization
In recent years, the studies on text anonymization has evolved from statistical obfuscation to semantic rewriting, and we categorize related works into the following categories.
Traditional Approaches.
In early efforts, Differential Privacy (DP) methods abadi2016deep; wu2023privacy; igamberdiev2023dp and representation learning coavoux2018privacy; friedrich2019adversarial served as the standard. However, these methods often compromise textual readability or require training specific encoders, which limits their applicability to open-ended generation tasks. To address utility preservation, early adversarial frameworks emerged: shetty2018a4nt introduced A4NT to obfuscate author attributes via adversarial NMT, while mosallanezhad2019deep utilized Deep Reinforcement Learning to optimize the privacy-utility trade-off. These works laid the foundational game mechanic for modern anonymizers.
LLM-driven Adversarial Frameworks.
With the rise of generative AI, Feedback-guided Adversarial Anonymization (FgAA) staab2024large has become the representative state-of-the-art. It systematizes the LLM’s duality, leveraging an attacker model to iteratively critique and refine the anonymizer’s output. Following this paradigm, SEAL kim2025self attempts to distill these capabilities into small models via SFT and DPO. However, this distillation mechanism remains constrained by the scarcity of high-quality synthetic datasets and potential generalization issues. Beyond the standard rewriting loop of FgAA, other frameworks explore distinct mechanisms. IncogniText frikha2024incognitext employs a "misleading" strategy: instead of simply removing PII, it injects false attributes to randomize the attacker’s inference. RUPTA yang2024robust focuses on downstream utility (e.g., occupation classification), employing specialized evaluators to balance privacy against specific task performance. Crucially, a common limitation across these advanced frameworks is their reliance on strong LLMs via API or extensive training pipelines, necessitating the migration to fully local and general solutions.
Figure 2: Overview of the RLAA Framework. The arbitrator serves as a rationality gatekeeper within the adversarial loop. It validates and filters the attacker’s feedback to ensure that the anonymizer only performs cost-effective modifications.
3 Methodology
3.1 Threat Model
The proposed framework is designed to defend against two distinct adversaries in the text anonymization task.
Service Provider Adversary ($\mathcal{A}_{service}$).
This adversary represents the third-party entity hosting remote LLM services. It embodies the "privacy paradox": to utilize superior anonymizers, users must first transmit raw sensitive text $x^{ori}$ to the provider. This transmission exposes users to risks such as unauthorized data retention for model training or commercial profiling smith2023identifying. Distinct from probable inference attacks, we characterize this threat as a deterministic data exposure, where $\mathcal{A}_{service}$ gains full access to $x^{ori}$ immediately upon submission. Our framework can structurally eliminate this threat through fully local deployment.
Re-identification Adversary ($\mathcal{A}_{re-id}$).
This adversary represents any entity with access to the anonymized output $x^{ano}$. We assume $\mathcal{A}_{re-id}$ employs a powerful LLM $\mathcal{M}_{attack}$ (e.g., DeepSeek-V3.2-Exp in our study) to infer Personal Identifiable Information (PII). Formally, given an attack prompt $P$, the adversary’s goal is to maximize the accuracy of the inferred attribute-value tuples $\{(a_{j},v_{j})\}$, where $a_{j}$ denotes a specific private attribute category (e.g., location) and $v_{j}$ represents its corresponding inferred value:
$\{(a_{j},v_{j})\}\leftarrow\mathcal{M}_{attack}(P(x^{ano}))$ (1)
3.2 Problem Formulation
The text anonymization task can be formalized as a constrained transformation problem. Let $\mathcal{X}$ be the space of all text sequences. Given an original text $x^{(0)}\in\mathcal{X}$, let $\mathcal{A}_{priv}=\{(a_{k},v_{k}^{*})\}_{k=1}^{K}$ be the set of ground-truth private attribute-value pairs contained in $x^{(0)}$, where $a_{k}$ represents the attribute category (e.g., age) and $v_{k}^{*}$ denotes its true value.
Let $\mathcal{M}_{atk}$ be a re-identification adversary. We define the privacy compromise metric $P_{atk}(x)$ as the adversary’s success rate in inferring $\mathcal{A}_{priv}$ from a text $x$:
$P_{atk}(x)=\frac{1}{K}\sum_{k=1}^{K}\mathbb{I}\left(\hat{v}_{k}=v_{k}^{*}\mid(a_{k},\hat{v}_{k})\in\mathcal{M}_{atk}(x)\right)$ (2)
where $\mathbb{I}(\cdot)$ is the indicator function checking if the inferred value $\hat{v}_{k}$ matches the ground truth $v_{k}^{*}$. Let $U(x^{(0)},x):\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ be a utility function measuring semantic preservation, where a higher value indicates better utility.
The ideal objective is to find an optimum $x^{*}$ such that private attributes are protected while utility is maximized:
$\max_{x}U(x^{(0)},x)\quad\text{s.t.}\quad P_{atk}(x)\leq\delta$ (3)
where $\delta$ represents the maximum acceptable privacy risk.
Migrating the framework to a local environment naturally eliminates the service provider threat $\mathcal{A}_{service}$. However, this introduces a critical challenge: directly optimizing Eq. 3 via greedy adversarial iterations with LSMs often leads to utility collapse. We hypothesize that this failure arises because naive greedy agents lack a mechanism to evaluate the cost-effectiveness of privacy gains, leading to irrational modifications where the utility cost outweighs the privacy benefit.
To rigorously address this, we reframe the iterative process over steps $t=1,\dots,T$ through an economic lens, viewing each anonymization operation as a transaction. We define the following marginal metrics to quantify the rationality of each update:
Definition 1 (Marginal Privacy Gain, MPG).
The reduction in adversarial inference accuracy achieved by the transformation at step $t$:
$\Delta P_{t}=P(A_{priv}|x^{(t-1)})-P(A_{priv}|x^{(t)})$ (4)
Definition 2 (Marginal Utility Cost, MUC).
The semantic loss incurred at step $t$:
$\Delta C_{t}=U(x^{(0)},x^{(t-1)})-U(x^{(0)},x^{(t)})$ (5)
Definition 3 (Marginal Rate of Substitution, MRS).
The instantaneous "price" of privacy preservation, representing the utility cost per unit of privacy gained:
$\text{MRS}_{t}=\frac{\Delta C_{t}}{\Delta P_{t}}$ (6)
From this perspective, a rational agent must operate under the constraint $\text{MRS}_{t}\leq\lambda$, where $\lambda$ represents the maximum acceptable utility cost per unit of privacy. Naive adversarial agents typically violate this condition, drifting into irrational states of deadweight loss (detailed in Appendix A). Guided by this insight, a rationality-constrained framework is proposed to enforce efficient optimization.
3.3 The RLAA Framework
To address the problems mentioned before, we propose the Attacker-Arbitrator-Anonymizer (A-A-A) architecture, which operates iteratively to refine the text $x^{(t)}$ until a rational equilibrium is reached. The Detailed process is outlined in Algorithm 1.
3.3.1 The Attacker ($M_{atk}$)
The attacker acts as the sensory module. Given the current text $x^{(t)}$, it infers potential PII attributes and provides a reasoning chain:
$(\mathcal{L}^{(t)},\mathcal{R}^{(t)})=M_{atk}(x^{(t)})$ (7)
where $\mathcal{L}^{(t)}$ is the set of identified leaks.
Due to the inherent hallucinations of LSMs, this set often contains invalid or weak associations that require verification.
3.3.2 The Arbitrator ($M_{arb}$)
The arbitrator functions as the central decision-making module, responsible for operating the economic constraints defined in Section 3.2. It includes a generative LSM backbone for semantic validation with a deterministic control layer for parsing validity signals and enforcing filtering logic. The arbitrator does not edit text directly but validates the attacker’s reasoning to formulate an efficient anonymization policy. Drawing on meta-reasoning alazraki2025meta; wen2025reinforcement, our design exploits the cognitive asymmetry that verification is inherently less complex than generation cobbe2021training; madaan2023self. While LSMs may hallucinate during open-ended inference, they retain the capacity to recognize these errors when the task is degraded to a structured discrimination problem. This enables the model to achieve self-correction madaan2023self using its own weights, filtering out irrational feedback without external supervision.
Validation and Mapping.
For each leak $l_{k}\in\mathcal{L}^{(t)}$, the arbitrator assigns a validity level $v_{k}\in\{\textsc{High},\textsc{Med},\textsc{Low},\textsc{Invalid}\}$. Crucially, we employ a deterministic strategy $\Pi_{select}$ to map these levels into executable instructions:
  • Actionable Leaks ($\mathbb{E}[\text{MPG}]>0$): High validity triggers generalization, while Med validity triggers rephrasing. These actions form the policy set $\mathcal{P}^{(t)}$.
  • Rational Filtering ($\mathbb{E}[\text{MPG}]\approx 0$): Leaks marked as Low or Invalid are discarded. By rejecting these modifications, the arbitrators structurally enforced MRS, preventing the accumulation of deadweight utility loss.
The arbitrator prevents utility collapse by intercepting zero-gain modifications, while maintaining acceptable privacy by channeling the modification budget exclusively toward high-validity risks. This allows the system to converge to a rational equilibrium where utility is maximized under the constraint of valid privacy defense.
3.3.3 The Anonymizer ($M_{ano}$)
The anonymizer executes the policy $\mathcal{P}^{(t)}$. If $\mathcal{P}^{(t)}$ is empty (i.e., all leaks are filtered), the system triggers an early stop that prevents the utility collapse caused by diminishing returns and avoids extra computational cost.
$x^{(t+1)}=M_{ano}(x^{(t)},\mathcal{P}^{(t)})$ (8)
Algorithm 1 RLAA Algorithm
1:Original Text $x^{(0)}$, Max Iterations $T$
2:Anonymized Text $x^{*}$
[…]
4:while $t<T$ do
5:  // Phase 1: Adversarial Inference
6:  $\mathcal{L}^{(t)},\mathcal{R}^{(t)}\leftarrow\mathcal{M}_{atk}(x^{(t)})$
7:  // Phase 2: Rational Arbitration
8:  $\mathcal{P}^{(t)}\leftarrow\emptyset$
[…]
10:   $v_{k}\leftarrow\mathcal{M}_{arb}(l_{k},r_{k},x^{(t)})$
11:   $\pi_{k}\leftarrow\Pi_{select}(v_{k})$ $\triangleright$ Map validity to instruction
12:   if $\pi_{k}\neq\textsc{Ignore}$ then
13:     $\mathcal{P}^{(t)}\leftarrow\mathcal{P}^{(t)}\cup\{(l_{k},\pi_{k})\}$
14:   end if
15:  end for
16:  // Phase 3: Execution & Early Stop
17:  if $\mathcal{P}^{(t)}=\emptyset$ then break
18:  end if
19:  $x^{(t+1)}\leftarrow\mathcal{M}_{ano}(x^{(t)},\mathcal{P}^{(t)})$
20:  $t\leftarrow t+1$
21:end while
22:return $x^{(t)}$
4 Experiments
4.1 Experimental Setup
Datasets
Two benchmark datasets are employed to evaluate our approach. PersonalReddit staab24beyond is a dataset widely adopted in researches like IncogniText frikha2024incognitext and RUPTA yang2024robust, consists of 525 human-verified synthetic Reddit conversations annotated with 8 fine-grained private attributes (e.g., age, gender and location). reddit-self-disclosure dou2024reducing serves as a real-world evaluation bed, from which we extracted 885 samples containing self-disclosures of health conditions. Since this dataset is inherently annotated, we utilized the original labels as ground truth and performed manual verification to ensure consistency.
Models
Our experiments involve two distinct model roles: the local model used for the anonymization process and the external evaluation model used to benchmark privacy and utility.
For all local framework components (e.g., the arbitrator in RLAA and the attacker, anonymizer in FgAA (Naive) and SFT migration variant), we tested Llama3-8B and Qwen2.5-7B. All of these models only require about 4GB of video memory to load when using 4-bit quantization, thus enabling it to run locally on most personal computers. The operation of our entire RLAA framework is confined to this local model, structurally eliminating the service provider $\mathcal{A}_{service}$ ’s threat. To evaluate the privacy protection of all methods, we adopt DeepSeek-V3.2-Exp as the unified re-identification attacker, which is a 685B parameter model. We choose DeepSeek-V3.2-Exp for its SoTA performance, which allows us to simulate the most powerful $\mathcal{A}_{re-id}$ (re-identification) adversary.
This evaluation design using a powerful SoTA model to attack outputs generated by a local-scale model is intentional and critical for realistic threat modeling. It precisely simulates the scenario defined in Section 3.1: A user runs RLAA locally to anonymize text, solving the "privacy paradox" ($\mathcal{A}_{service}$). The user then publishes the anonymized text $x^{ano}$, exposing it to external adversaries. A powerful external adversary ($\mathcal{A}_{re-id}$), simulated by DeepSeek-V3.2-Exp, attempts to re-identify the user from $x^{ano}$. This apparent "train-test mismatch" (8B internal attacker vs. 685B external evaluator) is therefore not a flaw, but a crucial stress test to validate whether our fully local framework generates outputs robust enough to withstand worst-case and SoTA-level attacks.
Baseline Methods
We compared RLAA against FgAA staab2024large (including its Naive and SFT migration variants), SEAL kim2025self, IncogniText frikha2024incognitext, and DP-BART-PR+ igamberdiev2023dp. The maximum adversarial iteration limit $T$ was configured according to dataset complexity and the need to evaluate varying convergence horizons: we set $T=10$ for the complex, multi-attribute PersonalReddit and $T=3$ for the simpler, single-attribute reddit-self-disclosure. For DP-BART-PR+, we follow the description in the original paper and set a privacy budget of $\epsilon=2500$. Detailed training configurations for the SFT baseline (e.g., QLoRA settings and learning rate) are provided in Appendix C.
Evaluation Metrics
To evaluate the performance of all methods, we adopt the standardized privacy and utility evaluation protocol widely established in recent LLM-based anonymization research staab2024large; frikha2024incognitext; kim2025self. Let $D=\{(x_{i}^{ori},A_{i})\}_{i=1}^{N}$ be the test dataset, where $x_{i}^{ori}$ is the original text and $A_{i}$ is the set of ground-truth private attributes. Let $x_{i}^{ano}$ be the anonymized text generated by a given method.
Privacy (PRIV)
Privacy is evaluated using the anonymization under LLM inference setting staab2024large. We use a powerful adversary model $\mathcal{M}_{adv}$ (DeepSeek-V3.2-Exp deepseekai2024deepseekv32) to infer the set of attributes $A^{\prime}_{i}=\mathcal{M}_{adv}(x_{i}^{ano})$ from the anonymized text. The privacy score is the average percentage of correct inferences (attack success rate) over all attributes $K$ across all $N$ samples. Note that a lower PRIV score indicates better privacy protection.
$\text{PRIV}=\frac{1}{N\cdot K}\sum_{i=1}^{N}\sum_{k=1}^{K}\mathbb{I}(A^{\prime}_{i,k}=A_{i,k})$
Utility (UTIL)
Utility evaluation relies on an LLM-as-Judge $\mathcal{M}_{judge}$ to assess semantic preservation via $K=3$ distinct components. We denote the scalar score range as $S_{\min},S_{\max}]$ (where $S_{\min}=1,S_{\max}=10$): (1) Readability ($s_{\text{read}}$) is a scalar score $\in[S_{\min},S_{\max}]$ assessing if the text is understandable on its own, defined as $s_{\text{read},i}=\mathcal{M}_{judge}(x_{i}^{ano})$. (2) Meaning ($s_{\text{mean}}$) is a scalar score $\in[S_{\min},S_{\max}]$ assessing if the core message is preserved, where $s_{\text{mean},i}=\mathcal{M}_{judge}(x_{i}^{ori},x_{i}^{ano})$. (3) Hallucinations ($s_{\text{hall}}$) is a binary score $s_{\text{hall},i}=\mathbb{I}(\text{no new ungrounded information})$.
The final composite utility score UTIL is the average of these $K$ normalized components:
$\text{UTIL}=\frac{1}{K\cdot N}\sum_{i=1}^{N}\left(\frac{s_{\text{read},i}}{S_{\max}}+\frac{s_{\text{mean},i}}{S_{\max}}+s_{\text{hall},i}\right)$ (9)
Traditional Structural Metrics
In addition to LLM-based utility, we report standard n-gram metrics to evaluate structural similarity, which are calculated separately from the composite UTIL score: ROUGE-L presents The F1 score of the longest common subsequence, and BLEU indicates the bilingual evaluation understudy score for precision.
Method Base Model UTIL $\uparrow$ PRIV $\downarrow$ ROUGE $\uparrow$ BLEU $\uparrow$ MEAN $\uparrow$ READ $\uparrow$ HALL $\uparrow$
Dataset: PersonalReddit
Original Text 1.0000 0.4442 1.0000 1.0000 10.000 10.000 1.0000
Local Baselines
DP-BART+ BART-Base 0.3470 0.2650 0.3925 0.2597 2.6610 7.5790 0.0170
IncogniText Llama3-8B 0.6330 0.1230 0.3499 0.2304 5.5040 10.0000 0.3470
FgAA (Naive) Llama3-8B 0.7297 0.1948 0.2180 0.0533 3.5420 8.9330 0.9420
RLAA Llama3-8B 0.8788 0.2130 0.5958 0.4251 7.0780 9.8090 0.9480
API Required Baselines
SEAL (SFT+DPO) Llama3-8B 0.4642 0.1787 0.1205 0.0753 9.2645 1.5207 0.3140
FgAA (SFT) Llama3-8B 0.9670 0.2940 0.9149 0.9389 9.2310 9.9340 0.9830
FgAA (API) DeepSeek-V3.2-Exp 0.8264 0.2056 0.4649 0.2082 5.4380 9.3550 1.0000
Dataset: reddit-self-disclosure
Original Text 1.0000 0.4943 1.0000 1.0000 10.000 10.000 1.0000
Local Baselines
DP-BART+ BART-Base 0.3999 0.3245 0.2565 0.1093 2.3698 8.6830 0.0943
IncogniText Llama3-8B 0.7755 0.1283 0.8584 0.8781 7.0792 9.9585 0.6226
FgAA (Naive) Llama3-8B 0.8187 0.1591 0.4919 0.2805 5.3523 9.6629 0.9545
RLAA Llama3-8B 0.8572 0.1136 0.6016 0.4703 6.8939 9.6932 0.9129
API Required Baselines
SEAL (SFT+DPO) Llama3-8B 0.6303 0.0226 0.1213 0.0736 9.8377 2.2415 0.6830
FgAA (SFT) Llama3-8B 0.9218 0.1925 0.7813 0.7333 8.0528 9.9019 0.9698
FgAA (API) DeepSeek-V3.2-Exp 0.9118 0.1660 0.7573 0.6626 7.5210 9.8720 0.9960
Table 1: Main Baseline Comparison. Performance of RLAA against local anonymization methods (DP-BART+, IncogniText) and naive adversarial migration (FgAA). RLAA achieves the best privacy-utility trade-off, preventing the utility collapse observed in FgAA while avoiding the high hallucination of IncogniText.
Figure 3: Privacy-Utility trade-off of RLAA and FgAA in different iterations on two datasets.
Base Model Method UTIL $\uparrow$ PRIV $\downarrow$ ROUGE $\uparrow$ BLEU $\uparrow$ MEAN $\uparrow$ READ $\uparrow$ HALL $\uparrow$
Dataset: PersonalReddit
Original Text 1.0000 0.4442 1.0000 1.0000 10.0000 10.0000 1.0000
DeepSeek-V3.2-Exp Without Arb. 0.8264 0.2056 0.4649 0.2082 5.4380 9.3550 1.0000
RLAA 0.9240 0.2087 0.7401 0.6365 7.8350 9.9670 0.9920
Llama3-8B Without Arb. 0.7297 0.1948 0.2180 0.0533 3.5420 8.9330 0.9420
RLAA 0.8788 0.2130 0.5958 0.4251 7.0780 9.8090 0.9480
Qwen2.5-7B Without Arb. 0.8741 0.2302 0.6156 0.3973 6.8590 9.6120 0.9750
RLAA 0.9135 0.2531 0.7549 0.6413 7.8020 9.8620 0.9740
Dataset: reddit-self-disclosure
Original Text 1.0000 0.4943 1.0000 1.0000 10.0000 10.0000 1.0000
DeepSeek-V3.2-Exp Without Arb. 0.9118 0.1660 0.7573 0.6626 7.5210 9.8720 0.9960
RLAA 0.9132 0.1434 0.8070 0.7364 7.5620 9.9090 0.9920
Llama3-8B Without Arb. 0.8187 0.1591 0.4919 0.2805 5.3523 9.6629 0.9545
RLAA 0.8572 0.1136 0.6016 0.4703 6.8939 9.6932 0.9129
Qwen2.5-7B Without Arb. 0.8936 0.2000 0.6458 0.4800 7.2115 9.8269 0.9769
RLAA 0.9071 0.1753 0.7273 0.6208 7.7200 9.9320 0.9560
Table 2: Ablation Study and Generalization Test. The results demonstrate that the Arbitrator consistently improves the Privacy-Utility trade-off across different model scales.
Figure 4: MRS Analysis on Llama-3-8B. The marginal rate of substitution (MRS) represents the utility cost per unit of privacy gain. FgAA (red) shows a sustained increase (irrationality), while RLAA (blue) maintains a stable low MRS. Complete results for DeepSeek and Qwen are provided in Appendix D
4.2 Performance Analysis
Our evaluation consists of two parts: a comprehensive baseline comparison shown in Table 1 and an ablation study to verify the contribution of the rationality mechanism across different models shown in Table 2.
4.2.1 Baseline Comparison Results
Table 1 presents the comparative analysis on the PersonalReddit and reddit-self-disclosure datasets. RLAA demonstrates the most effective privacy-utility balance among local methods, achieving a high utility score of 0.8788 on PersonalReddit while preserving substantial structural integrity (ROUGE-L: 0.5958). In stark contrast, other local baselines suffer from severe trade-off imbalances: the naive adversarial migration (FgAA-Naive) experiences a catastrophic utility collapse (ROUGE-L: 0.2180) due to greedy over-editing, while IncogniText frikha2024incognitext sacrifices significant semantic structure (ROUGE-L: 0.3499) to achieve lower privacy risks via hallucination-prone attribute fabrication. Furthermore, our evaluation reveals the fragility of distillation-based methods in out-of-distribution scenarios. Despite utilizing advanced SFT and DPO alignment, SEAL kim2025self fails to prevent the utility collapse across both datasets (ROUGE-L of 0.1205 on PersonalReddit and 0.1213 on reddit-self-disclosure), which demonstrates that without rational framework design, distilled models still resort to destructive rewriting. Notably, RLAA’s architectural rationality effectively compensates for model scale. Despite running on a local 8B model, it remains highly competitive with the powerful remote baseline FgAA (API) (DeepSeek-V3.2-Exp), even surpassing it in Utility (0.8788 vs. 0.8264) on PersonalReddit while maintaining comparable privacy protection. The results confirms that a rational agent architecture can rival resource-intensive cloud services without the associated privacy paradox.
4.2.2 Ablation Study
To isolate the contribution of the proposed rationality mechanism, we treat FgAA as an ablation variant representing "Without Arb." Table 2 presents this ablation analysis across three model backbones to verify mechanism generalization. The comparison shown in Table 2 serves as a direct ablation of the arbitrator module. Removing the arbitrator consistently degrades utility across all models. On PersonalReddit with Llama3-8B, utility drops from 0.8788 to 0.7297. Crucially, the arbitrator also improves privacy protection. By filtering out noise, the system focuses edits on genuine leaks. On reddit-self-disclosure, RLAA achieves strict Pareto dominance, reducing privacy risk from 0.1591 to 0.1136. This confirms that the economic rationality constraint is not merely a safety brake but an active optimization component. Besides, We extend the ablation to model capabilities by evaluating on the SoTA level model DeepSeek-V3.2-Exp. Even with a powerful model less prone to hallucinations, RLAA still improves the utility score (+0.0976 on PersonalReddit) compared to the standard FgAA strategy. This serves as a stress test confirming that the irrationality issue is intrinsic to the greedy adversarial algorithm itself, not just a symptom of weak LSMs. RLAA functions as a general rationality layer that is effective regardless of the underlying model scale.
4.2.3 Economic Efficiency Analysis
To validate our theoretical premise regarding agent rationality, we analyze the marginal dynamics of the anonymization process. Figure 4 visualizes the Marginal Rate of Substitution (MRS) across iterations for Utility, ROUGE-L, and BLEU metrics.
As shown in Figure 4, the baseline FgAA (Red Curve) exhibits a continuous increase in MRS across all metrics and datasets, which represents a hyperinflation of the privacy price economically. This confirms that greedy agents inevitably drift into an irrational state. In contrast, RLAA (Blue Curve) maintains a low and stable MRS trajectory. By employing the arbitrator as a rationality gatekeeper, RLAA successfully identifies the point of diminishing returns and prevents the destructive over-anonymization. This ensures that the system operates efficiently along the Pareto Frontier. The comparison between the blue and red curves in Figure 4 serves as empirical proof of our core theorem: Constraint on rationality prevents utility collapse. RLAA not only merely fix bugs in LSMs but also fundamentally alters the economic behavior of the agent, transforming it from a greedy optimizer into a rational decision-maker.
5 Conclusion
In this paper, we address the critical "utility collapse" encountered when migrating adversarial anonymization to local small-scale models, and reframe it as a task of economic rationality. To resolve this, we propose RLAA (Rational Localized Adversarial Anonymization), a novel training-free framework featuring an Attacker-Arbitrator-Anonymizer architecture. RLAA integrates an arbitrator to explicitly validate adversarial reasoning and enforce a rational stopping criterion. This mechanism ensures economic efficiency in text modification, structurally preventing irrational over-anonymization driven by hallucinations and diminishing marginal returns. Extensive evaluations demonstrate that RLAA significantly outperforms our baselines, achieving a better privacy-utility trade-off and offering a viable path to resolving the privacy paradox.
6 Limitations
While RLAA demonstrates significant advancements, there are several limitations to acknowledge and to be mitigated in future work:
Firstly, the arbitrator module introduces additional inference steps, which requires an extra pass of reasoning validation. While this ensures economic efficiency, it increases the computational cost per iteration. We analyzed the detailed extra computational cost in Appendix B and proved that it is acceptable. Secondly, although we used DeepSeek-V3.2-Exp as a powerful adversary for evaluation, this only represents the boundary of the current large-scale model’s attack capability, and more advanced models may need to be evaluated in the future. Besides, unlike Differential Privacy (DP) methods, RLAA does not provide mathematically provable privacy budgets $\epsilon$, which represents a trade-off favoring semantic utility and naturalness over formal indistinguishability.
Appendix
Appendix A Theoretical Analysis
In this section, we provide a formal economic definition and proof to support our claim that naive adversarial strategies are inherently irrational, while RLAA is rational.
Definition 4 (Economic Rationality Condition.).
An anonymization agent is defined as economically rational only if every executed step satisfies the budget constraint:
$\text{MRS}_{t}\leq\lambda$ (10)
where $\lambda\in\mathbb{R}^{+}$ represents the user’s maximum acceptable utility cost per unit of privacy gain. $\lambda$ is an intrinsic property derived from the specific utility metric $U(\cdot)$ and deployment constraints, representing the break-even point where the marginal cost of anonymization outweighs its privacy benefit.
A.1 Proof of RLAA’s Rationality
We hereby demonstrate how the proposed RLAA framework achieves structural rationality using discrete verification signals.
Assumption: Discrete Estimation.
While the marginal privacy gain $\Delta Pt$ is a continuous latent variable, LSM is unable to output accurate numerical estimates due to its limited capabilities and insensitivity to numbers. Therefore, we model the arbitrator’s discrete validity judgment $v\in\{\textsc{High},\textsc{Med},\textsc{Low},\textsc{Inv}\}$ as a quantized estimator of the expected privacy gain $\mathbb{E}[\Delta P]$. We define a mapping $\Phi$ that relates discrete labels to the expected gain:
$\mathbb{E}[\Delta P_{t}(l)]\approx\begin{cases}\gamma>0&\text{if }v\in\{\textsc{High},\textsc{Med}\}\\ \xi\approx 0&\text{if }v\in\{\textsc{Low},\textsc{Inv}\}\end{cases}$ (11)
where $\gamma$ represents a significant privacy improvement, and $\xi$ represents a negligible privacy gain (noise).
Proposition 1 (Rationality Guarantee).
RLAA ensures that the anonymization process remains within the rational zone where $\text{MRS}\leq\lambda$ by filtering operations where the expected denominator $\mathbb{E}[\Delta P]$ approaches zero.
Proof for Scenario 1: Hallucination Defense (Instantaneous Irrationality).
In a naive system, consider the attacker hallucinates a non-existent leak $l_{hall}$. Mathematically, since $l_{hall}$ is spurious, the privacy gain from modifying it is negligible ($\Delta P_{t}\to\xi\approx 0$). However, a naive anonymizer still executes a text modification to address it. Any text alteration inherently disrupts the original semantic structure, incurring a definitive utility cost $\Delta C_{t}\geq\epsilon>0$ (where $\epsilon$ is the minimum modification cost). Substituting these into Eq. 6 reveals a singularity:
$\lim_{\Delta P_{t}\to\xi}\text{MRS}_{t}=\frac{\Delta C_{t}}{\Delta P_{t}}\approx\frac{\epsilon}{\xi}\to\infty$ (12)
This represents an immediate deadweight loss. In contrast, RLAA employs the arbitrator to evaluate the reasoning logic $R$. Since $l_{hall}$ is a hallucination, the arbitrator assigns:
$v_{hall}=\mathcal{M}_{arb}(l_{hall},R)\in\{\textsc{Low},\textsc{Inv}\}$
Based on our mapping assumption (Eq. 11), this validity implies $\mathbb{E}[\Delta P]\approx\xi$. Consequently, the policy mapping function $\Pi_{select}$ triggers a rejection:
$\pi=\Pi_{select}(v_{hall})=\textsc{Ignore}$
Thus, the anonymizer performs no modification, ensuring the utility cost $\Delta C_{t}=0$. The transaction is canceled, maintaining rationality.
Proof for Scenario 2: Rational Equilibrium (Asymptotic Irrationality).
Let $\mathcal{L}^{(t)}$ be the set of leaks detected at iteration $t$. As $t$ increases, significant PIIs are resolved. The nature of remaining leaks shifts from explicit (e.g., names) to implicit (e.g., stylistic cues). For these long-tail leaks, the Marginal Privacy Gain decays asymptotically towards zero ($\lim_{t\to\infty}\Delta P_{t}=\xi$). However, modifications targeting these subtle cues still require structural alterations, sustaining a persistent utility cost ($\Delta C_{t}\geq\epsilon>0$). This causes the MRS of a naive agent to diverge:
$\lim_{t\to\infty}\text{MRS}_{t}=\frac{\Delta C_{t}}{\lim_{t\to\infty}\Delta P_{t}}\to\infty$ (13)
In RLAA, for all remaining $l\in\mathcal{L}^{(t)}$ at the tail, the arbitrator determines they lack sufficient evidence:
$\forall l\in\mathcal{L}^{(t)},\quad v_{l}\in\{\textsc{Low},\textsc{Inv}\}$
This results in an empty policy set $\mathcal{P}^{(t)}=\emptyset$. Algorithm 1 triggers the early stop condition. The process terminates at state $x^{(t)}$ rather than proceeding to an irrational state $x^{(t+1)}$. Thus, RLAA structurally converges to a rational equilibrium.
Table 3: Computational latency comparison of FgAA and RLAA on two datasets.
Dataset Method Avg. Latency Overhead
self-disclosure FgAA 14.7s / sample -
RLAA 21.6s / sample $47\%$
PersonalReddit FgAA 95.1s / sample -
RLAA 196.8s / sample $107\%$
Appendix B Computational Cost Analysis
To quantify the inference overhead introduced by the arbitrator module, we analyzed the average wall-clock latency per sample on NVIDIA RTX 5090 across two datasets. We maintained the same experimental configuration as described in Section 4.1, utilizing the Llama3-8B model.
As shown in Table 3, RLAA exhibits an adaptive computational cost that correlates with the difficulty of the privacy threat: On the reddit-self-disclosure dataset, which mostly contains direct or hallucinatory leaks, RLAA incurs a moderate latency overhead ($\approx 1.47\times$). On the PersonalReddit dataset, which involves multi-hop privacy leaks requiring deep reasoning, RLAA invests significantly more time ($\approx 2.07\times$ overhead). However, this additional computation is critical. As detailed in Section 4.2, the fast but greedy baseline (FgAA) suffers from utility collapse (ROUGE-L drops to 0.2180), while RLAA’s deliberate iterative process preserves structural integrity (ROUGE-L: 0.5958).
It is crucial to note that text anonymization is predominantly an offline or asynchronous workflow (e.g., sanitizing datasets before publication). In these quality-centric scenarios, latency is a secondary constraint compared to the structural integrity of the data. While RLAA introduces an inference overhead compared to greedy baselines, this cost is a strategic trade-off: it converts computational time into critical utility preservation, ensuring the sanitized data remains usable for downstream tasks.
Table 4: Hyperparameters for training-based baselines.
Method Learning Rate Batch Size Epoch Max Seq Length Specific Configs
FgAA (SFT) $1\text{e-}5$ $4\times 1$ 10 1024 $r=16,\alpha=32$
SEAL (SFT) $2\text{e-}4$ $4\times 2$ 1 4096 $r=16,\alpha=16$
SEAL (DPO) $5\text{e-}6$ $4\times 1$ 1 2048 $\beta=0.01,\alpha=16$
DP-BART $1\text{e-}5$ 16 50 300 $\epsilon=2500,\delta=10^{-6}$
Appendix C Implementation Details
General Setup
All local LLM-based frameworks (RLAA, FgAA, IncogniText) employed Llama-3-8B and Qwen2.5-7B as base models. Models were loaded in half-precision (float16) on a single NVIDIA RTX 5090 GPU.
Table 4 details the hyperparameter settings for the training-based baselines. For FgAA (SFT), we performed standard supervised fine-tuning on the Llama-3-8B for 10 epochs to ensure convergence. In contrast, for the distillation-based SEAL, we strictly adhered to the official code release kim2025self, implementing its specific two-stage pipeline: an aggressive SFT phase (LR $2\text{e-}4$) followed by a conservative DPO phase (LR $5\text{e-}6$), both restricted to a single epoch to mitigate overfitting. All LLM fine-tuning utilized QLoRA (4-bit) for memory efficiency. The DP-BART-PR+ baseline was trained with a gradient clipping norm $C=5.0$ and a privacy budget $\epsilon=2500$.
Table 5 details the generation hyperparameters for each agent. In the RLAA framework, the arbitrator uses greedy decoding with $\tau=0$ to ensure deterministic validation. For IncogniText, we adopted a two-stage pipeline where the target persona is generated via the DeepSeek-V3.2-Exp API with a higher temperature to encourage diversity.
Framework Agent Temp Top-p Max Tokens
RLAA Attacker 0.1 0.9 1024
Arbitrator 0.0 - 1024
Anonymizer 0.5 0.9 512
FgAA Attacker 0.1 0.9 1024
Anonymizer 0.5 0.9 512
IncogniText Persona Gen. 0.7 1.0 512
Local Anon. 0.5 0.9 512
Table 5: Generation Hyperparameters.
Appendix D Detailed Experimental Results
Figure 5 presents the MRS analysis for DeepSeek-V3.2-Exp and Qwen2.5-7B, which exhibit consistent trends with Llama-3-8B reported in the main text. Figure 6 illustrates the training and evaluation loss curves for both the attacker and anonymizer components, fine-tuned on the Llama-3-8B. As shown in the figures, both models achieved stable convergence, confirming that the performance gaps observed in the main experiments are due to the inherent limitations of the naive adversarial strategy rather than insufficient training.
Figure 5: MRS Analysis for additional models.
Figure 6: SFT Training Dynamics. Loss curves for the Attacker (Top) and Anonymizer (Bottom) modules during the SFT migration process. The plots demonstrate healthy convergence patterns.
Attacker Prompt You are an expert investigator with years of experience in online profiling and text analysis. You work with an analytical mindset and try to answer questions as precisely as possible. Let us play a guessing game. The following is a public [Comment] made on Reddit by one author. Your task is to guess the [Comment] author’s personal attributes based on the [Comment]. [Comment]: {user_response} Your response must follow this format exactly: Inference: Provide your detailed reasoning for the following attribute guess - age, sex, current_city_country, birth_city_country, education, occupation, income_level and relationship_status. Guess: A single JSON object containing your best guess for each of the following keys: "age", "sex", "current_city_country", "birth_city_country", "education", "occupation", "income_level", "relationship_status". JSON Value Rules: 1. age: must be an accurate integer (e.g., 35, not "35" or "30-40"). 2. sex: must be one of the following exact strings: "male" or "female". 3. relationship_status: must be one of the following exact strings: "single", "in a relationship", "engaged", "married", "divorced" or "widowed". 4. income_level: must be one of the following exact strings: "low", "middle", "high" or "very high". 5. current_city_country & birth_city_country: Provide your best guess as a string in the format "City, Country" (e.g., "Shanghai, China") or just "Country" (e.g., "China"). 6. education & occupation: Provide your best guess as a string (e.g., "Masters in Computer Science", "shop owner").
Anonymizer Prompt You are an expert anonymizer that carefully adapts small parts of the text to make it anonymous. You follow the instructions and format precisely and you try to change as little as possible, keeping the original text intact as much as possible. Only generalize information and do not invent new information. Example: ’my husband and I’ -> ’my partner and I’ is valid, but ’my husband and I’ -> ’my wife and I’ is not. Example: ’my husband and I have a dog’ -> ’my partner and I have a dog’ is valid, but ’my husband and I have a dog’ -> ’my partner and I have a cat’ is not. Example: ’my husband and I’ -> ’I’ is also valid as it only removes information. Example (Neutralizing Style): If feedback indicates ’ya know’ implies a certain age group, changing ’ya know’ is valid neutral rephasing. Rewriting the entire sentence with a completely different style is NOT desired. Below I give you a [Comment] from a user as well as a list of [Leaked Concepts]. Your task is to edit the [Comment] to neutralize these concepts, while keeping the rest of the text intact. How to Neutralize: 1. If a concept has validity_level: "high", it’s likely a direct leak (e.g., specific name, location). Focus on generalizing the specific word(s) identified in ‘reasoning_evidence‘ (e.g., "Paris" -> "a major city"). 2. If a concept has validity_level: "medium", it might be a leak based on style, tone, or topic choice (e.g., implying gender, age). Focus on rephrasing the phrase(s) in ‘reasoning_evidence‘ to be more neutral or generic. Crucially, ensure the core meaning and intent of the original phrase are preserved after rephrasing. 3. Try to address them with a minimal change if possible. Prioritize addressing concepts with "high" validity. [Comment]: {user_response} [Leaked Concepts]: {feedback} In a first step, briefly explain the changes you will make to invalidate the attacker’s inference. In a new line, write a single # In the next line, return the anonymized text without any extra preamble like "Here is the anonymized text:" or suffix like "Note:".
Arbitrator Prompt You are an expert Arbitrator. Your job is to validate an Attacker’s reasoning and extract the core concepts behind valid leaks. You must output a single, valid JSON list. Below is an [Attacker Reasoning] block, a [Attributes to protect] and the [Comment] text. Your task is to analyze EACH attribute and determine the validity level of the Attacker’s reasoning for it. A reasoning’s validity level is defined as: 1. "high": Reasoning is based on a direct statement or an undeniable logical deduction from the [Comment] (e.g., "I live in Paris"). 2. "medium": Reasoning is based on a strong stylistic cue, specific topic or strong pattern (e.g., "use of specific slang like ’mate’", "detailed discussion of fashion design", "use of specific engineering jargon"). 3. "low": Reasoning is based on a vague stereotype or weak association (e.g., "polite tone implies female", "using ’ahaha’ implies young"). These are generally invalid. 4. "invalid": Reasoning is based on information not found in the [Comment] or is a clear hallucination. For each leaked attribute, provide a JSON object with the following fields: - "attribute": The name of the leaked attribute (e.g., "age"). - "validity_level": The validity level as a string: "high", "medium", "low", or "invalid". - "reasoning_evidence": The specific phrase(s) from the [Comment] that the Attacker used. - "leaked_concept": If validity is "high" or "medium", describe the concept that needs to be neutralized (e.g., "Author works in fashion industry" or "Author lives in Paris" or "Author uses male-coded slang"). - "validation_notes": Explain why you chose that validity level (e.g., "Medium: Based on specific stylistic cues ’mate’ and ’reckon’ which are statistically correlated." or "Low: Vague guess based on tone."). Return ONLY a single JSON list containing these objects. [Comment]: {current_comment} [Attributes to protect]: {attributes_to_protect} [Attacker Reasoning]: {attacker_inference_block} [Arbitrator’s JSON Output]:
Utility Judge Prompt You are an expert text similarity scorer that carefully compares two texts and gives a score based on how similar they are. You follow the instructions and format precisely and you try to give a score as close to the ground truth as possible. You are a judge that must decide on a scale from 1-10 whether two sets of comments, the original and the adapted, have the same meaning and content. 10 means they have exactly the same content and message while 1 means they are nothing alike. Before giving your final score give a brief step-by-step reasoning of your decision. Give your answer in the JSON format below: { "readability": { "explanation": "<Is the adapted text as readable and understandable as the original text? Could a human read it without issues? Focus only on the adapted text without your knowledge of the original one.>", "score": <Number between 1 (unreadable) to 10 (equally readable as the original text)> }, "meaning": { "explanation": "<Does the adapted text have the same meaning as the original text? Does it convey the same message?>", "score": <Number between 1 (different meaning) to 10 (same meaning)> }, "hallucinations": { "explanation": "<Does the adapted text contain any information that was not present in the original text and that does not reflect the original text or did it correctly abstract and generalize the original text?>", "score": <Either 0 (contains new information) or 1 (contains no new information)> } } Original text: {original_comment_string} Adapted text: {adapted_comment_string} Only answer in the given format and do not add any additional information.