### v2

# Look Twice before You Leap: A Rational Agent Framework for Localized Adversarial Anonymization

Abstract
Current LLM-based text anonymization frameworks usually rely on remote API services from powerful LLMs, which creates an inherent privacy paradox: users must disclose data to untrusted third parties for guaranteed privacy preservation. Moreover, directly migrating current solutions to local small-scale models (LSMs) offers a suboptimal solution with severe utility collapse. Our work argues that this failure stems not merely from the capability deficits of LSMs, but significantly from the inherent irrationality of the greedy adversarial strategies employed by current state-of-the-art (SOTA) methods. To address this, we propose Rational Localized Adversarial Anonymization (RLAA), a fully localized and training-free framework featuring an Attacker-Arbitrator-Anonymizer architecture. We model the anonymization process as a trade-off between Marginal Privacy Gain (MPG) and Marginal Utility Cost (MUC), and demonstrate that greedy strategies tend to drift into an irrational state. Instead, RLAA introduces an arbitrator that acts as a rationality gatekeeper, validating the attacker’s inference to filter out feedback providing negligible privacy benefits. This mechanism promotes a rational early-stopping criterion, and structurally prevents utility collapse. Extensive experiments on different benchmarks demonstrate that RLAA achieves a superior privacy-utility trade-off compared to strong baselines.
Look Twice before You Leap: A Rational Agent Framework for
Localized Adversarial Anonymization
Donghang Duan1, Xu Zheng1, Yuefeng He1, Chong Mu1, Leyi Cai2, Lizong Zhang1 1School of Computer Science and Engineering, University of Electronic Science and Technology of China 2School of Information, Renmin University of China
1 Introduction
Large language models (LLMs) are increasingly deployed to process real-world text containing sensitive Personal Identifiable Information (PII), such as medical records, legal documents and online self-disclosures bommasani2021opportunities; weidinger2021ethical; deusser2025survey. To comply with stringent regulatory requirements like GDPR and CCPA albrecht2016gdpr; bonta2022california, effective text anonymization has become a prerequisite for the responsible use of such data. However, this situation presents a tricky trade-off between semantic utility and privacy, where over-anonymization destroys the semantic value while under-anonymization invites re-identification risks, which is particularly acute in sensitive domains like smart healthcare and legal technology im2024exploring; morris2025diri; liu2025legal. The rapid evolution of LLMs has exacerbated this situation because of their dual roles as context-aware anonymizers and superior attackers wang2025unique. As a result, traditional NER-based anonymization methods are increasingly inadequate dernoncourt2017identification, motivating a shift toward semantic anonymization that conceals latent identity cues beyond surface-level entities loiseau2025tau.
To address this situation, advanced anonymization frameworks based on LLMs have emerged. As the current state-of-the-art (SOTA) framework, Feedback-guided Adversarial Anonymization (FgAA) staab2024large typically employs a dynamic adversarial game: an attacker model attempts to re-identify personal information from the text generated by an anonymizer model, and the anonymizer uses the inference feedback to instruct the anonymization process for refinement. Despite its strong empirical performance, FgAA heavily relies on the capabilities of powerful or closed-source LLMs like GPT-4 achiam2023gpt and DeepSeek-V3.2-Exp deepseekai2024deepseekv32, which are typically accessible only via third-party APIs. This architectural dependency gives rise to a fundamental privacy paradox: to anonymize sensitive data, users must first transmit the raw content to external and uncontrollable service providers, which constitutes a critical privacy vulnerability smith2023identifying and renders the architecture unacceptable in high-sensitivity scenarios feretzakis2024trustworthy. Although subsequent works like IncogniText frikha2024incognitext and SEAL kim2025self aim to mitigate this risk through localized deployment, they compromise semantic utility by injecting ungrounded hallucinations or require complex training pipelines.
Figure 1: Utility Collapse of FgAA’s Naive Migration.
A natural and simple solution to the privacy paradox is to migrate adversarial anonymization frameworks to fully localized environments using local small-scale models (LSMs). However, our empirical investigation reveals that such a naive migration leads to severe utility collapse. As illustrated in Figure 1, the naive migration on Llama3-8B causes destructive over-editing. It indiscriminately strips away non-sensitive context and stylistic nuances, reducing the rich original narrative to a generic and vacuous summary. Besides, it is observed that this collapse persists even when implementing instruction refinement (e.g., outputting "unknown" for inferences with low confidence). This observation demonstrates that heuristic constraints alone cannot rectify the inherent irrationality of the greedy decision-making process.
We argue that this failure cannot be attributed solely to the limited capabilities of LSMs, but instead arises from how greedy adversarial strategies behave under imperfect inferences. From an economic perspective, the anonymization can be viewed as a sequence of decisions balancing Marginal Privacy Gain (MPG) against Marginal Utility Cost (MUC) mas1995microeconomic. By quantifying this ratio as the Marginal Rate of Substitution ($\text{MRS}=\frac{\text{MUC}}{\text{MPG}}$), the primary driver of the utility collapse becomes conceptually clear through this economic lens: Current greedy strategies tend to drift into an economically inefficient state due to LSMs’ hallucinations and the law of diminishing returns, where the model executes destructive edits for non-existent or negligible leaks (see theoretical analysis in Appendix A). In this case, MUC remains positive while MPG approaches zero, pushing MRS to inefficient levels. One concurrent study SEAL kim2025self aims to distill knowledge from a powerful LLM to an LSM through Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) rafailov2023direct to avoid disclosing raw data to third-party APIs. However, our empirical analysis in Section 4.3 shows that it still fails to handle the rationality of anonymization. Besides, SEAL requires additional resources for post-training while high-quality synthetic datasets remain scarce and difficult to construct effectively yukhymenko2024synthetic.
To overcome both the irrationality of naive migration and the data dependency of distillation-based approaches, we propose RLAA (Rational Localized Adversarial Anonymization) [^1] Our code and data are available at https://github.com/SowingG2333/RLAA., a training-free framework designed for fully localized deployment. The core idea of RLAA is to constrain the anonymization process via architectural rationality, thereby compensating for the capability deficits of LSMs without relying on explicit numerical control or parameter fine-tuning. RLAA introduces a novel Attacker–Arbitrator–Anonymizer (A–A–A) architecture, in which the arbitrator acts as a rational gatekeeper between attack feedback and anonymization actions. Moving beyond the paradigm of blindly leaping into modifications, RLAA compels the arbitrator to scrutinize the attacker’s inferences through a validation mechanism inspired by meta-reasoning alazraki2025meta; wen2025reinforcement. This mechanism structurally prevents utility collapse by rejecting destructive edits driven by invalid feedback, and maintains robust privacy protection by focusing the modification budget solely on genuine leaks. Crucially, this design leverages the cognitive asymmetry that verification is less complex and more reliable than generation for LSMs, which may hallucinate during open-ended inference but retain the ability to recognize errors in structured discrimination tasks cobbe2021training; wang2022self; madaan2023self; guan2024language. We conducted a comprehensive evaluation including baseline comparisons, ablation studies and generalization stress tests across several mainstream LLMs (Llama3-8B llama3modelcard, Qwen2.5-7B qwen2.5 and DeepSeek-V3.2-Exp) on the PersonalReddit and reddit-self-disclosure datasets. Additionally, we validated the agent’s rational decision-making process through a quantitative economic efficiency analysis based on the MRS metric. The results empirically demonstrate that RLAA structurally prevents utility collapse and achieves a superior privacy-utility trade-off compared to strong baselines. In summary, the main contributions of this study are as follows:
  • We argue that migrating adversarial anonymization to local environments is crucial for eliminating API-based privacy risks, while a naive migration to LSMs results in utility collapse. The collapse in local adversarial anonymization is identified as a symptom of the economic irrationality inherent in greedy strategies, rather than merely capability deficits.
  • RLAA is proposed as a localized and training-free framework featuring an Attacker-Arbitrator-Anonymizer (A-A-A) architecture. By introducing an arbitrator, it structurally promotes rational decision-making to prevent utility collapse.
  • Extensive experiments demonstrate that RLAA achieves a superior privacy-utility trade-off compared to multiple strong baselines. It achieves Pareto dominance on reddit-self-disclosure and demonstrate its mechanism generalization across different base models.
2 Related Work
2.1 Inference Attacks through LLMs
With increased reasoning capabilities, LLMs have evolved into potent privacy attackers capable of automatically inferring personal attributes from unstructured text wang2025unique. staab24beyond demonstrated that models can accurately deduce fine-grained PII from casual online comments. Extending to high-stakes domains, nyffenegger2024anonymity revealed that LLMs could re-identify individuals in court decisions. Recent audits by panda2025daiq further confirmed that this inference capability is consistent across various model architectures. These studies show that LLMs significantly lower the threshold for privacy violations, enabling re-identification attacks to occur on an unprecedented scale staab24beyond.
2.2 Text Anonymization
Recent advancements in text anonymization have marked a fundamental evolution from early statistical obfuscation techniques toward sophisticated semantic rewriting mechanisms. Generally, these studies can be divided into two categories:
Traditional Approaches. In early efforts, Differential Privacy (DP) methods abadi2016deep; wu2023privacy; igamberdiev2023dp and representation learning coavoux2018privacy; friedrich2019adversarial served as the standard. However, these methods often compromise textual readability or require training specific encoders, which limits their applicability to open-ended generation tasks. To address utility preservation, shetty2018a4nt introduced A4NT to obfuscate author attributes, while mosallanezhad2019deep utilized deep reinforcement learning to optimize the privacy-utility trade-off. These works laid the foundational mechanism for modern anonymizers.
LLM-driven Adversarial Frameworks. With the rise of generative AI, Feedback-guided Adversarial Anonymization (FgAA) staab2024large established the prevailing paradigm for state-of-the-art anonymization. It systematizes the LLM’s duality, leveraging an attacker model to iteratively critique and refine the anonymizer’s output. Following this paradigm, SEAL kim2025self attempts to distill these capabilities into smaller models through SFT and DPO from multi-round teacher trajectories. However, this mechanism remains constrained by synthetic data scarcity yukhymenko2024synthetic and potential generalization issues. IncogniText frikha2024incognitext employs a misleading strategy, injecting false attributes to randomize attacker inferences rather than simply removing PII. Beyond general-purpose anonymization, RUPTA yang2025robust addresses a complementary problem: optimizing anonymized text for specific downstream tasks (e.g., occupation classification) using specialized task evaluators. While orthogonal to our focus on general semantic preservation, RUPTA shares the common reliance on API-based LLMs. Crucially, a common limitation across these advanced frameworks is their reliance on strong LLMs via API or extensive training pipelines. Moreover, methods following the greedy adversarial paradigm lack rational decision-making mechanisms, making them prone to utility collapse when deployed with smaller local models.
Figure 2: The RLAA Framework. Utilizing an Attacker-Arbitrator-Anonymizer architecture, the arbitrator acts as a rationality gatekeeper. It validates attacker inferences to filter out ghost leaks with negligible privacy benefits, structurally preventing utility collapse caused by irrational greedy strategies.
3 Methodology
3.1 Threat Model
RLAA is designed to defend against two distinct adversaries in the text anonymization task:
Semi-honest Service Provider ($\mathcal{A}_{serv}$). This adversary represents a third-party entity that hosts remote LLM services. It causes the privacy paradox: to utilize superior anonymizers, users must first transmit raw sensitive text $x^{ori}$ to the provider. This transmission exposes users to risks such as unauthorized data retention for model training or commercial profiling smith2023identifying. Distinct from probable inference attacks, we characterize this threat as a deterministic data exposure, where $\mathcal{A}_{serv}$ gains full access to $x^{ori}$ immediately upon submission.
Re-identification Adversary ($\mathcal{A}_{re-id}$). This adversary represents any entity with access to the anonymized output $x^{ano}$. We assume $\mathcal{A}_{re-id}$ employs a powerful LLM $\mathcal{M}_{atk}$ (e.g., DeepSeek-V3.2-Exp in our study) to infer PII. Given an attack prompt $P$, the adversary’s goal is to maximize the accuracy of the inferred attribute-value tuples $\{(a_{j},v_{j})\}$, where $a_{j}$ denotes a specific private attribute category (e.g., location) and $v_{j}$ represents its corresponding inferred value (e.g., "Paris, France"):
$\{(a_{j},v_{j})\}\leftarrow\mathcal{M}_{atk}(P(x^{ano}))$ (1)
3.2 Problem Formulation
The text anonymization task can be formalized as a constrained transformation problem. Let $\mathcal{X}$ be the space of all text sequences. Given an original text $x^{(0)}\in\mathcal{X}$, let $\mathcal{A}_{priv}=\{(a_{k},v_{k}^{*})\}_{k=1}^{K}$ be the set of "private attribute-true value" pairs contained in $x^{(0)}$, where $a_{k}$ represents the attribute category (e.g., age) and $v_{k}^{*}$ denotes its true value.
Let $\mathcal{M}_{atk}$ be a re-identification adversary. We define the privacy compromise metric $P_{atk}(x)$ as the adversary’s success rate in inferring $\mathcal{A}_{priv}$ from a text $x$:
$P_{atk}(x)=\frac{1}{K}\sum_{k=1}^{K}\mathbb{I}\left(\hat{v}_{k}=v_{k}^{*}\mid(a_{k},\hat{v}_{k})\in\mathcal{M}_{atk}(P(x))\right)$ (2)
where $\mathbb{I}(\cdot)$ is the indicator function checking if the inferred value $\hat{v}_{k}$ matches the ground truth $v_{k}^{*}$.
Simultaneously, let $U(x^{(0)},x):\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ be a utility function measuring semantic preservation, where a higher value indicates better utility. The ideal objective is to find an optimum $x^{*}$ such that private attributes are protected while utility is maximized:
$\max\quad U(x^{(0)},x)\quad\text{s.t.}\quad P_{atk}(x)\leq\delta$ (3)
where $\delta$ represents the upper bound of acceptable privacy risk.
Migrating the framework to a local environment naturally eliminates the service provider threat $\mathcal{A}_{serv}$. However, achieving the objective in Eq. 3 is non-trivial. Existing adversarial frameworks typically approximate the optimum $x^{*}$ via an iterative sequence $x^{(0)}\to x^{(1)}\dots\to x^{(T)}$, where each step attempts to reduce privacy risk. Our empirical investigation reveals a critical challenge: directly applying greedy adversarial iterations with LSMs in this process often leads to utility collapse. We hypothesize that this failure arises because naive greedy agents lack a mechanism to evaluate the cost-effectiveness of privacy gains. To formally address this, we reframe this iterative process over steps $t=1,2\dots T$ through an economic lens, viewing each anonymization operation as a transaction. The following marginal metrics are defined to quantify the rationality of each update.
Definition 1 (Marginal Privacy Gain, MPG).
The reduction in adversarial inference accuracy achieved by the transformation at step $t$:
$\Delta P_{t}=P_{atk}(x^{(t-1)})-P_{atk}(x^{(t)})$ (4)
where $x^{(t-1)}$ represents the intermediate text generated at the previous iteration.
Definition 2 (Marginal Utility Cost, MUC).
The semantic loss incurred at step $t$:
$\Delta C_{t}=U(x^{(0)},x^{(t-1)})-U(x^{(0)},x^{(t)})$ (5)
Definition 3 (Marginal Rate of Substitution, MRS).
The instantaneous price of privacy preservation, representing the utility cost per unit of privacy gained:
$\text{MRS}_{t}=\frac{\Delta C_{t}}{\Delta P_{t}}$ (6)
From this perspective, a rational agent should ideally align with the principle that $\text{MRS}_{t}\leq\lambda$, where $\lambda$ represents the maximum rational utility cost per unit of privacy. Naive greedy agents inherently tends to violate this condition and drift into irrational states of deadweight loss where $\text{MRS}_{t}\to\infty$. Guided by this insight, RLAA is proposed to implicitly enforce this budget constraint $\lambda$ through architectural rationality, thereby structurally preventing utility collapse. Detailed theoretical analysis is shown in Appendix A.
3.3 The RLAA Framework
To address the problems mentioned before, we propose the Attacker-Arbitrator-Anonymizer (A-A-A) architecture, which operates iteratively to refine the text $x^{(t)}$. For a detailed algorithmic procedure, please refer to the Algorithm 1 in Appendix C.
3.3.1 The Attacker ($M_{atk}$)
The attacker acts as the sensory module. Given the current text $x^{(t)}$, it infers potential PII attributes and provides a reasoning chain:
$(\mathcal{L}^{(t)},\mathcal{R}^{(t)})=M_{atk}(x^{(t)})$ (7)
where $\mathcal{L}^{(t)}$ is the set of identified leaks and $\mathcal{R}^{(t)}$ is the set of corresponding inferences.
3.3.2 The Arbitrator ($M_{arb}$)
The arbitrator functions as the central control interface that regulates anonymization decisions considering the rationality constraints defined in Section 3.2. It includes a generative LSM backbone for semantic validation with a deterministic control layer for parsing validity signals and enforcing filtering logic. Instead of explicitly optimizing marginal trade-offs, which is unreliable given LSM’s limited sensitivity to fine-grained scalar signals sun2025numerical, the arbitrator validates attacker inferences by constraining the LSM to a structured discrimination task. By leveraging the greater reliability of verification over generation guan2024language, this design enables self-correction and structurally blocks irrational feedback from driving utility collapse.
For each leak $l_{k}\in\mathcal{L}^{(t)}$ observed by the attacker, the arbitrator assigns a validity level $v_{k}\in\{\textsc{High},\textsc{Med},\textsc{Low},\textsc{Invalid}\}$. From the economic perspective defined in Section 3.2, this validity level serves as a discrete estimator of the MPG, distinguishing valid risks ($\Delta P_{t}>0$) from negligible noise ($\Delta P_{t}\approx 0$). Accordingly, we define the selection policy $\Pi_{select}$ as a mapping from validity levels to anonymization mandates:
$\Pi_{select}(v_{k})=\begin{cases}\textsc{Execute}&\text{if }v_{k}\in\{\textsc{High},\textsc{Med}\}\\ \textsc{Ignore}&\text{if }v_{k}\in\{\textsc{Low},\textsc{Invalid}\}\end{cases}$ (8)
The Execute branch captures leakage with significant returns, ensuring a finite MRS. Conversely, the Ignore branch filters out zero-gain noise where $\Delta P_{t}\to 0$. By rejecting these transactions, the arbitrator prevents the system from drifting into the MRS singularity identified in Eq. 6, thereby structurally prevents utility collapse.
3.3.3 The Anonymizer ($M_{ano}$)
The anonymizer executes the refined policy $\mathcal{P}^{(t)}$, which aggregates the actionable leaks validated by the arbitrator:
$\mathcal{P}^{(t)}=\{(l_{k},\pi_{k})\mid l_{k}\in\mathcal{L}^{(t)},\Pi_{select}(v_{k})=\textsc{EXECUTE}\}$ (9)
Based on this set, the text update is governed by:
$x^{(t+1)}=\begin{cases}M_{ano}(x^{(t)},\mathcal{P}^{(t)})&\text{if }\mathcal{P}^{(t)}\neq\emptyset\\ x^{(t)}&\text{if }\mathcal{P}^{(t)}=\emptyset\end{cases}$ (10)
When $\mathcal{P}^{(t)}$ is empty (i.e., all leaks are filtered by the arbitrator), the system triggers the early stop branch. This mathematically forces the iteration to converge to the fixed point $x^{(t+1)}=x^{(t)}$, thereby structurally preventing the utility collapse caused by diminishing returns.
4 Experiments
4.1 Experimental Setup
Datasets. Two benchmark datasets are employed to evaluate our approach. PersonalReddit staab24beyond is a dataset widely adopted in researches like IncogniText and RUPTA, consists of 525 human-verified synthetic Reddit conversations annotated with 8 fine-grained private attributes (e.g., age, gender and location). reddit-self-disclosure dou2024reducing serves as a real-world evaluation bed, from which 885 samples containing self-disclosures of health conditions are extracted. Since this dataset is inherently annotated, we utilize the original labels as ground truth and perform manual verification to ensure consistency.
Models. Our experiments involve two distinct model roles: a local model for the anonymization process and an external evaluation model for benchmarking privacy and utility. We employ Llama3-8B and Qwen2.5-7B as local backbones. With 4-bit quantization, both models require approximately 4GB VRAM, enabling fully consumer-grade local deployment and eliminating reliance on external services under the $\mathcal{A}_{serv}$ threat model. For evaluation, we utilize the 685B DeepSeek-V3.2-Exp to instantiate a strong $\mathcal{A}_{re-id}$ adversary aligned with the scenario defined in Section 3.1. Unlike proprietary APIs (e.g., GPT-4), which impose strict cost and access constraints, DeepSeek enables large-scale and strong attacks at negligible cost due to its open-source nature. This makes it particularly suitable for modeling an economically rational adversary operating under realistic resource constraints.
Baseline Methods. We compared the performance of RLAA against FgAA staab2024large (including its Naive and SFT migration variants), SEAL kim2025self, IncogniText frikha2024incognitext, and DP-BART-PR+ igamberdiev2023dp. The maximum adversarial iteration limit $T$ was configured according to dataset complexity and the need to evaluate varying convergence horizons: we set $T=10$ for the multi-attribute PersonalReddit and $T=3$ for the single-attribute reddit-self-disclosure. Detailed training configurations are provided in Appendix C.
4.2 Evaluation Metrics
To evaluate the performance of all methods, we adopt the standardized privacy and utility evaluation protocols widely established in recent LLM-based anonymization research staab2024large; frikha2024incognitext; kim2025self. Let $D=\{(x_{i}^{ori},A_{i})\}_{i=1}^{N}$ be the test dataset, where $x_{i}^{ori}$ is the original text and $A_{i}$ is the set of ground-truth private attributes. Let $x_{i}^{ano}$ be the anonymized text generated by a given method.
Privacy (PRIV). The performance on privacy preservation is evaluated using the anonymization under LLM inference setting staab2024large. We use a powerful adversary model $\mathcal{M}_{adv}$ (DeepSeek-V3.2-Exp) to infer the set of attributes $A^{\prime}_{i}=\mathcal{M}_{adv}(x_{i}^{ano})$ from the anonymized text. The privacy score is the average attack success rate over all attributes $K$ across all $N$ samples. A lower PRIV Score indicates better privacy protection.
$\text{PRIV}=\frac{1}{N\cdot K}\sum_{i=1}^{N}\sum_{k=1}^{K}\mathbb{I}(A^{\prime}_{i,k}=A_{i,k})$ (11)
Method Base Model UTIL $\uparrow$ PRIV $\downarrow$ ROUGE $\uparrow$ BLEU $\uparrow$ MEAN $\uparrow$ READ $\uparrow$ HALL $\uparrow$
PersonalReddit
Original Text - 1.0000 0.4442 1.0000 1.0000 10.000 10.000 1.0000
Local Methods DP-BART+ BART-Base 0.3470 0.2650 0.3925 0.2597 2.6610 7.5790 0.0170
IncogniText Llama3-8B 0.6330 0.1230 0.3499 0.2304 5.5040 10.0000 0.3470
FgAA-Naive Llama3-8B 0.7297 0.1948 0.2180 0.0533 3.5420 8.9330 0.9420
RLAA Llama3-8B 0.8788 0.2130 0.5958 0.4251 7.0780 9.8090 0.9480
API Required SEAL Llama3-8B 0.4642 0.1787 0.1205 0.0753 9.2645 1.5207 0.3140
FgAA-SFT Llama3-8B 0.9670 0.2940 0.9149 0.9389 9.2310 9.9340 0.9830
FgAA-API DeepSeek-V3.2-Exp 0.8264 0.2056 0.4649 0.2082 5.4380 9.3550 1.0000
reddit-self-disclosure
Original Text - 1.0000 0.4943 1.0000 1.0000 10.000 10.000 1.0000
Local Methods DP-BART+ BART-Base 0.3999 0.3245 0.2565 0.1093 2.3698 8.6830 0.0943
IncogniText Llama3-8B 0.7755 0.1283 0.8584 0.8781 7.0792 9.9585 0.6226
FgAA-Naive Llama3-8B 0.8187 0.1591 0.4919 0.2805 5.3523 9.6629 0.9545
RLAA Llama3-8B 0.8572 0.1136 0.6016 0.4703 6.8939 9.6932 0.9129
API Required SEAL Llama3-8B 0.6303 0.0226 0.1213 0.0736 9.8377 2.2415 0.6830
FgAA-SFT Llama3-8B 0.9218 0.1925 0.7813 0.7333 8.0528 9.9019 0.9698
FgAA-API DeepSeek-V3.2-Exp 0.9118 0.1660 0.7573 0.6626 7.5210 9.8720 0.9960
Table 1: Main Baseline Comparison. Performance of RLAA against local anonymization baselines (DP-BART+, IncogniText and FgAA-Naive) and API required baselines (SEAL, FgAA-SFT and FgAA-API).
Figure 3: Privacy-Utility Trade-off. RLAA achieves superior trade-offs compared to FgAA across iterations on two datasets. The trade-off dynamics for structural metrics (ROUGE-L/BLEU) are detailed in Appendix D.
Base Model Method UTIL $\uparrow$ PRIV $\downarrow$ ROUGE $\uparrow$ BLEU $\uparrow$ MEAN $\uparrow$ READ $\uparrow$ HALL $\uparrow$
PersonalReddit
Original Text - 1.0000 0.4442 1.0000 1.0000 10.0000 10.0000 1.0000
DeepSeek-V3.2-Exp w/o Arb. 0.8264 0.2056 0.4649 0.2082 5.4380 9.3550 1.0000
RLAA 0.9240 0.2087 0.7401 0.6365 7.8350 9.9670 0.9920
Llama3-8B w/o Arb. 0.7297 0.1948 0.2180 0.0533 3.5420 8.9330 0.9420
RLAA 0.8788 0.2130 0.5958 0.4251 7.0780 9.8090 0.9480
Qwen2.5-7B w/o Arb. 0.8741 0.2302 0.6156 0.3973 6.8590 9.6120 0.9750
RLAA 0.9135 0.2531 0.7549 0.6413 7.8020 9.8620 0.9740
reddit-self-disclosure
Original Text - 1.0000 0.4943 1.0000 1.0000 10.0000 10.0000 1.0000
DeepSeek-V3.2-Exp w/o Arb. 0.9118 0.1660 0.7573 0.6626 7.5210 9.8720 0.9960
RLAA 0.9132 0.1434 0.8070 0.7364 7.5620 9.9090 0.9920
Llama3-8B w/o Arb. 0.8187 0.1591 0.4919 0.2805 5.3523 9.6629 0.9545
RLAA 0.8572 0.1136 0.6016 0.4703 6.8939 9.6932 0.9129
Qwen2.5-7B w/o Arb. 0.8936 0.2000 0.6458 0.4800 7.2115 9.8269 0.9769
RLAA 0.9071 0.1753 0.7273 0.6208 7.7200 9.9320 0.9560
Table 2: Ablation and Generalization. Comparison between the full RLAA framework and the baseline without arbitrator. The arbitrator consistently improves utility across different model scales and datasets.
Figure 4: Cumulative MRS Analysis of Llama3-8B. The figure displays Llama3-8B’s cumulative MRS during the anonymization process on two datasets. FgAA (Red) shows a sustained increase, whereas RLAA (Blue) maintains a stable low MRS. The remaining results for DeepSeek-V3.2-Exp and Qwen2.5-7B are provided in Appendix D.
Utility (UTIL). The utility evaluation relies on an LLM-as-Judge $\mathcal{M}_{judge}$ (DeepSeek-V3.2-Exp) zhang2024small to assess semantic preservation via $K$ distinct components (where $K=3$). We denote the scalar score range as $S_{\min},S_{\max}]$ (where $S_{\min}=1,S_{\max}=10$): (1) Readability ($s_{\text{read}}$) is a scalar score $\in[S_{\min},S_{\max}]$ assessing if the text is understandable on its own, defined as $s_{\text{read},i}=\mathcal{M}_{judge}(x_{i}^{ano})$. (2) Meaning ($s_{\text{mean}}$) is a scalar score $\in[S_{\min},S_{\max}]$ assessing if the core message is preserved, where $s_{\text{mean},i}=\mathcal{M}_{judge}(x_{i}^{ori},x_{i}^{ano})$. (3) Hallucinations ($s_{\text{hall}}$) is a binary score $s_{\text{hall},i}=\mathbb{I}(\text{no new ungrounded information})$. And the final composite utility score UTIL is the average of these $K$ normalized components:
$\text{UTIL}=\frac{1}{K\cdot N}\sum_{i=1}^{N}\left(\frac{s_{\text{read},i}}{S_{\max}}+\frac{s_{\text{mean},i}}{S_{\max}}+s_{\text{hall},i}\right)$ (12)
Traditional Structural Metrics. In addition to LLM-based UTIL Score, we report standard n-gram metrics to evaluate structural similarity: ROUGE denotes the ROUGE-L F1 Score (lin2004rouge) based on the longest common subsequence, and BLEU (papineni2002bleu) indicates the bilingual evaluation understudy score for precision.
4.3 Performance Analysis
Our analysis consists of a comprehensive baseline comparison shown in Table 1 and an ablation study to verify the contribution of the rationality mechanism across different models detailed in Table 2.
4.3.1 Baseline Comparison Results
Table 1 presents the comparative analysis on the PersonalReddit and reddit-self-disclosure datasets. RLAA achieves the optimal privacy-utility balance among local methods (UTIL=0.8788/0.8572). In contrast, FgAA-Naive experiences a pronounced utility collapse due to greedy over-editing, which is further confirmed by our human evaluation in Appendix E, where RLAA achieves a dominant 88.4% win rate over FgAA-Naive. Besides, IncogniText sacrifices significant semantic structure (HALL=0.3470) by fabricating attributes that introduce ungrounded and potentially misleading contents. Distillation-based methods reveal a critical failure mode. SEAL achieves strong privacy (PRIV=0.18/0.02) but catastrophic utility loss (UTIL=0.46/0.63). We argue this stems from distilling multi-round teacher trajectories that lack early-stopping signals: the student inherits privacy-seeking aggressiveness but never learns when anonymization should stop. Conversely, our FgAA-SFT variant (detailed in Appendix C) corrects for over-caution (PRIV=0.29/0.19). This reveals distillation’s fundamental limitation: it transfers editing behaviors but not the meta-judgment of rational stopping. In contrast, RLAA’s training-free arbitrator structurally enforces this rationality, rivaling FgAA-API (DeepSeek-685B) with 8B models.
4.3.2 Ablation Study
To isolate the arbitrator’s contribution, we treat FgAA as the ablation baseline without arbitrator w/o Arb.. Table 2 shows consistent utility gains across all models. Crucially, RLAA achieves Pareto dominance on reddit-self-disclosure, improving both privacy (0.1591→0.1136) and utility (0.8187→0.8572), which demonstrates that rationality constraints actively optimize rather than merely constrain. Even the powerful DeepSeek-685B benefits from the arbitrator (+0.0976 utility on PersonalReddit), confirming irrationality stems from the greedy algorithm itself, not model weakness. This validates RLAA as a general architectural principle: structural rationality compensates for algorithmic myopia regardless of model scale.
4.3.3 Economic Efficiency Analysis
To validate our theoretical premise regarding agent rationality, we analyze the marginal dynamics of the anonymization process. Figure 4 visualizes the MRS across iterations for Utility Score. FgAA (Red Curve) exhibits a continuous increase in MRS across all metrics and datasets, which confirms that the greedy strategy tends to drift into an economically inefficient state. In contrast, RLAA (Blue Curve) maintains a low and stable MRS trajectory, showing economically rational anonymization behaviors. The comparison between the blue and red curves serves as an empirical proof of our hypothesis: Constraint on rationality prevents utility collapse. Besides, further quantitative analysis in Appendix D reveals a counter-intuitive capability-rationality paradox: the SOTA level model DeepSeek-V3.2-Exp exhibits a steeper irrationality drift than the smaller model Qwen2.5-7B. This observation suggests that scaling capabilities alone cannot resolve rationality alignment failures.
5 Conclusion
We reframe the utility collapse in local adversarial anonymization as an economic rationality problem and propose RLAA, a training-free framework with an Attacker-Arbitrator-Anonymizer architecture. By validating inferences and enforcing rational stopping, RLAA structurally prevents over-anonymization from hallucinations and diminishing returns. Extensive evaluations demonstrate that RLAA significantly outperforms our baselines, achieving a better privacy-utility trade-off and offering a viable path to resolving the privacy paradox.
Limitations
While RLAA effectively addresses the utility collapse in local anonymization, we acknowledge several limitations:
Inference Latency and the Path to Internalization. The introduction of the arbitrator module inevitably incurs additional inference costs due to the extra validation pass. Although our analysis in Appendix B indicates that this overhead remains acceptable for offline sanitization, it poses challenges for real-time applications. However, we view the current RLAA framework as a crucial inference-time alignment stage. Future work can leverage RLAA to generate high-quality and rationality-aligned trajectories. By fine-tuning models on these trajectories, we can internalize the arbitrator’s external constraints directly into the model’s parameters (training-time alignment), thereby eliminating the runtime overhead of an external gatekeeper.
Evolving Adversarial Capabilities. Our evaluation relies on DeepSeek-V3.2-Exp to simulate a powerful adversary. While this represents a current SOTA threat model, the landscape of LLM attacks is rapidly evolving. RLAA provides empirical defense, but we cannot theoretically guarantee immunity against future models with significantly stronger reasoning capabilities or attackers possessing extensive auxiliary background knowledge.
Lack of Provable Guarantees. Unlike Differential Privacy (DP) mechanisms, RLAA does not provide mathematically provable privacy budgets ($\epsilon$). This is a known trade-off in LLM-based text anonymization: strictly enforcing DP often results in significant degradation of semantic readability and utility. Our approach prioritizes semantic preservation and empirical safety, positioning it as a pragmatic solution rather than a mathematically guaranteed one.
Ethics Statement
We acknowledge the dual-use risks of LLMs and employ our adversarial framework strictly for defensive evaluation. Our approach enhances privacy by enabling fully localized deployment, thereby eliminating data exposure to third-party APIs. We clarify that RLAA provides empirical defense against SOTA adversaries, rather than provable guarantees like Differential Privacy. Regarding data usage, we rely exclusively on established public datasets (PersonalReddit and reddit-self-disclosure) and strictly adhere to ethical standards by avoiding any new collection of private data.
Appendix
Appendix A Theoretical Analysis
In this section, we provide a formal economic definition and analytical derivation showing that naive adversarial strategies are inherently irrational, while RLAA achieves structural rationality.
Definition 4 (Economic Rationality Condition).
An anonymization agent is defined as economically rational only if every executed step satisfies the budget constraint:
$\text{MRS}_{t}\leq\lambda$ (13)
where $\lambda\in\mathbb{R}^{+}$ represents the user’s maximum acceptable utility cost per unit of privacy gain. $\lambda$ is an intrinsic property derived from the specific utility metric $U(\cdot)$ and deployment constraints, representing the break-even point where the marginal cost of anonymization outweighs its privacy benefit.
Assumption 1 (Discrete Estimation).
We model the arbitrator’s discrete validity judgment $v\in\{\textsc{High},\textsc{Med},\textsc{Low},\textsc{Inv}\}$ as a quantized estimator of the expected privacy gain $\mathbb{E}[\Delta P]$. We define a mapping that relates discrete labels to the expected gain:
$\mathbb{E}[\Delta P_{t}(l)]\approx\begin{cases}\gamma>0&\text{if }v\in\{\textsc{High},\textsc{Med}\}\\ \xi\approx 0&\text{if }v\in\{\textsc{Low},\textsc{Inv}\}\end{cases}$ (14)
where $\gamma$ is a significant privacy improvement, and $\xi$ is negligible noise.
Proposition 1 (Implicit Budget Enforcement).
The arbitrator’s discrete mechanism creates a binary decision boundary that is functionally equivalent to an economic agent operating with an implicit budget $\lambda$.
Derivation.
Assuming the utility cost for any atomic edit is lower-bounded by $\Delta C_{t}\geq\epsilon$, we define the MRS lower bound for actionable leaks and the MRS upper bound for noise:
$\text{MRS}_{valid}\approx\frac{\epsilon}{\gamma},\quad\text{MRS}_{noise}\approx\frac{\epsilon}{\xi}\to\infty$ (15)
Under the assumption of correct estimation, there exists a significant separation gap: $\text{MRS}_{valid}\ll\text{MRS}_{noise}$. Therefore, the Arbitrator’s policy $\Pi_{select}$ effectively implements a budget constraint $\lambda$ located within this gap:
$\exists\lambda\in\left[\frac{\epsilon}{\gamma},\frac{\epsilon}{\xi}\right)\text{ s.t. }\forall v,\ \Pi_{select}(v)\neq\textsc{Ignore}\iff\text{MRS}(v)\leq\lambda$ (16)
This mechanism ensures the system structurally prioritizes high-return transactions while rejecting deadweight losses. ∎
Corollary 1 (Hallucination Defense - Instantaneous Rationality).
In a naive system, hallucinated leaks ($l_{hall}$) incur definitive utility costs ($\epsilon$) for negligible privacy gains ($\xi$). We can formally derive the irrationality of executing such edits as a singularity in the MRS:
$\lim_{\Delta P\to\xi}\text{MRS}(l_{hall})\approx\frac{\epsilon}{\xi}\to\infty$ (17)
By applying Proposition 1, RLAA identifies such instances as Low/Invalid validity. Since the implicit MRS exceeds any rational budget $\lambda$, the policy triggers a rejection $\Pi_{select}=\textsc{Ignore}$. Thus, the transaction cost is forced to zero ($\Delta C_{t}=0$), thereby avoiding the immediate deadweight loss.
Corollary 2 (Rational Equilibrium - Asymptotic Rationality).
As the iteration $t$ increases, the process enters a long-tail phase where remaining leaks are negligible with $\lim_{t\to\infty}\Delta P_{t}=\xi$. A greedy agent fails to stop because it lacks a mechanism to evaluate the diverging cost-benefit ratio:
$\lim_{t\to\infty}\text{MRS}_{greedy}=\frac{\Delta C_{t}}{\lim_{t\to\infty}\Delta P_{t}}\to\infty$ (18)
In contrast, RLAA applies Proposition 1 to the entire set of remaining candidates $\mathcal{L}^{(t)}$. Since $\forall l\in\mathcal{L}^{(t)},\text{MRS}(l)>\lambda$, the policy set becomes empty ($\mathcal{P}^{(t)}=\emptyset$). This explicitly triggers the algorithmic early stop, causing the system to converge to a stable rational equilibrium state $x^{(t)}$:
$x^{(t+1)}=x^{(t)}\quad(\text{Stop Condition})$ (19)
This mechanism structurally counteracts the drift into utility collapse observed in naive baselines.
Dataset Method Avg. Latency Overhead
self-disclosure FgAA 14.7s / sample -
RLAA 21.6s / sample $47\%$
PersonalReddit FgAA 95.1s / sample -
RLAA 196.8s / sample $107\%$
Table 3: Quantitative analysis of inference latency. Comparison between FgAA and RLAA using Llama3-8B on an NVIDIA RTX 5090 GPU.
Method Learning Rate Batch Size Epoch Max Length Specific Configs
FgAA (SFT) $1\text{e-}5$ $4\times 1$ 10 1024 $r=16,\alpha=32$
SEAL (SFT) $2\text{e-}4$ $4\times 2$ 1 4096 $r=16,\alpha=16$
SEAL (DPO) $5\text{e-}6$ $4\times 1$ 1 2048 $\beta=0.01,\alpha=16$
DP-BART $1\text{e-}5$ 16 50 300 $\epsilon=2500,\delta=10^{-6}$
Table 4: Training hyperparameters for Baselines. All utilize 4-bit QLoRA to ensure memory efficiency.
Appendix B Inference Latency Analysis
To quantify the inference overhead introduced by the arbitrator module, we analyzed the average wall-clock latency per sample on NVIDIA RTX 5090 across two datasets. The evaluation follows the same experimental configuration as described in Section 4.1, utilizing the Llama3-8B model. As shown in Table 3, RLAA exhibits an adaptive inference latency that correlates with the difficulty of the privacy threat: On the reddit-self-disclosure dataset, which mostly contains direct or hallucinatory leaks, RLAA incurs a moderate latency overhead ($\approx 1.47\times$). On the PersonalReddit dataset, which involves multi-hop privacy leaks requiring deep reasoning, RLAA invests significantly more time ($\approx 2.07\times$). However, this additional computation is critical. As detailed in Section 4.3, the fast but greedy baseline (FgAA) suffers from utility collapse (ROUGE-L→0.2180), while RLAA’s deliberate iterative process preserves structural integrity (ROUGE-L=0.5958). It is crucial to note that text anonymization is predominantly an offline or asynchronous workflow (e.g., sanitizing datasets before publication). In these quality-centric scenarios, latency is a secondary constraint compared to the structural integrity of the data. While RLAA introduces an inference overhead compared to greedy baselines, this cost is a strategic trade-off: it converts computational time into critical utility preservation, ensuring the text remains semantic utility.
Algorithm 1 Rational Anonymization
1:Original Text $x^{(0)}$, Max Iterations $T$
2:Anonymized Text $x^{*}$
[…]
4:while $t<T$ do
5:  // Phase 1: Adversarial Inference
6:  $\mathcal{L}^{(t)},\mathcal{R}^{(t)}\leftarrow\mathcal{M}_{atk}(x^{(t)})$
7:  // Phase 2: Rational Arbitration
8:  $\mathcal{P}^{(t)}\leftarrow\emptyset$
[…]
10:   $v_{k}\leftarrow\mathcal{M}_{arb}(l_{k},r_{k},x^{(t)})$
11:   $\pi_{k}\leftarrow\Pi_{select}(v_{k})$
12:   if $\pi_{k}\neq\textsc{Ignore}$ then
13:     $\mathcal{P}^{(t)}\leftarrow\mathcal{P}^{(t)}\cup\{(l_{k},\pi_{k})\}$
14:   end if
15:  end for
16:  // Phase 3: Execution & Early Stop
17:  if $\mathcal{P}^{(t)}=\emptyset$ then break
18:  end if
19:  $x^{(t+1)}\leftarrow\mathcal{M}_{ano}(x^{(t)},\mathcal{P}^{(t)})$
20:  $t\leftarrow t+1$
21:end while
22:return $x^{(t)}$
Appendix C Implementation Details
This appendix provides implementation configurations, including the algorithmic procedure in Algorithm 1, computational environment, training recipes and generation hyperparameters.
Base Models & Environment. All local LLM-based frameworks (RLAA, FgAA, IncogniText) employed Llama-3-8B and Qwen2.5-7B as base models. To align with consumer-grade deployment scenarios, all models were loaded in half-precision (float16) on a single NVIDIA RTX 5090 GPU.
Training Configurations for Baselines. Table 4 details the specific hyperparameter settings for all training-based baselines.
  • FgAA (SFT): We performed standard supervised fine-tuning on Llama-3-8B for 10 epochs to ensure convergence. This variant serves as an experimental ablation to probe whether training alone can impart rational anonymization behavior. Specifically, the teacher model generates one anonymization trajectory for each instance and is required to output "unknown" when an attribute cannot be reasonably inferred, aiming to produce a more rational student. All LLM fine-tuning utilized QLoRA (4-bit) for memory efficiency.
  • SEAL: For this distillation-based baseline, we strictly adhered to the official implementation, including its two-stage pipeline: an SFT phase followed by a conservative DPO phase, executed as specified by the released recipe.
  • DP-BART-PR+: According to the official code, this baseline was trained with a gradient clipping norm $C=5.0$ and a privacy budget $\epsilon=2500$.
Framework Agent Temp Top-p Max Tokens
RLAA Attacker 0.1 0.9 1024
Arbitrator 0.0 - 1024
Anonymizer 0.5 0.9 512
FgAA Attacker 0.1 0.9 1024
Anonymizer 0.5 0.9 512
IncogniText Persona Gen. 0.7 1.0 512
Local Anon. 0.5 0.9 512
Table 5: Specific Generation Hyperparameters. We use greedy decoding for the arbitrator to ensure deterministic validation and nucleus sampling for the anonymizer to maintain output diversity.
Generation Hyperparameters. Table 5 details the generation hyperparameters for each agent.
  • RLAA Settings: The arbitrator uses a temperature of 0 to ensure deterministic validity judgments, while the attacker and anonymizer use a slight temperature (0.1 and 0.5) to balance creativity and instruction following.
  • IncogniText Pipeline: We adopt a two-stage pipeline where the target persona is pre-generated offline via DeepSeek-V3.2-Exp to encourage stylistic diversity. Crucially, API usage is strictly limited to this one-time data preparation phase, ensuring the subsequent anonymization process remains fully localized.
Figure 5: Privacy-utility trade-offs via structural metrics (ROUGE-L and BLEU). Results are shown for PersonalReddit (Left) and reddit-self-disclosure (Right), demonstrating RLAA’s resistance to structural collapse.
Appendix D Detailed Experimental Results
This section expands on the main analysis by providing structural evaluations, training stability checks and a quantitative economic analysis that reveals a counter-intuitive alignment paradox.
Structural Privacy-Utility Trade-offs. Figure 5 plots the trade-off dynamics for structural metrics (ROUGE-L and BLEU). Consistent with the composite Utility Score, the FgAA-Naive baseline suffers from structural collapse as it aggressively purges information. In contrast, RLAA maintains high structural integrity along the Pareto frontier. This confirms that the arbitrator effectively distinguishes between necessary privacy edits and destructive structural damage.
Figure 6: Cumulative MRS Profiles across Different Model Scales. RLAA consistently reduces the MRS while revealing the capability-rationality paradox where stronger models exhibit higher over-editing tendencies in greedy baselines.
Figure 7: Extended MRS Dynamics. These results across Utility, ROUGE-L, and BLEU metrics confirm the generalization of RLAA’s rational decision-making.
The Capability-Rationality Paradox: A Quantitative Analysis. We extend the economic analysis to Qwen2.5-7B and DeepSeek-V3.2-Exp, summarizing the comparative rationality profiles in Figure 6 and detailing their iterative dynamics in Figure 7. To quantify the impact of RLAA across different model capabilities, we calculate the Rationality Gain (percentage reduction of MRS) in Table 6. This combined assessment reveals a distinct capability-rationality paradox: Despite being the SOTA level model, DeepSeek-V3.2-Exp exhibits the highest rationality gain. As visualized in Figure 6 and quantified in Table 6, RLAA reduces DeepSeek’s MRS by a massive 66.9%, indicating that without RLAA, it functions as the least economically rational agent due to safety over-alignment. In contrast, Qwen2.5-7B shows the least rationality gain of 28.9%. Its greedy baseline (FgAA’s MRS of 1.80) is naturally closer to the rational equilibrium, suggesting a more balanced inherent alignment for anonymization tasks. These results confirm that RLAA acts as an adaptive rationality gatekeeper, ensuring a consistent rational equilibrium ($\text{MRS}\approx 1.1\text{-}1.7$) regardless of the base model’s inherent bias.
Metric DeepSeek Llama Qwen
FgAA’s MRS 3.32 3.80 1.80
RLAA’s MRS 1.10 1.74 1.28
Rationality Gain 66.9% 54.2% 28.9%
Table 6: Quantifying Rationality Correction. RLAA demonstrates significant efficiency improvements over FgAA, particularly for DeepSeek-V3.2-Exp.
Training Convergence Verification. To ensure the baseline failure is not due to under-fitting, Figure 8 illustrates the SFT training dynamics. Both Attacker and Anonymizer modules exhibit healthy convergence with stable loss reduction. This rules out optimization failure and confirms that the performance gap stems from the inherent economic irrationality of the greedy strategy itself.
Figure 8: SFT Training Dynamics. The stable loss reduction and smooth convergence observed across Attacker (Left) and Anonymizer (Right) rule out optimization failure or under-fitting as the underlying cause for baseline utility collapse.
Appendix E Human Evaluation
To empirically validate the superiority of our rationality mechanism and confirm the usability of the generated text, we conducted a three-way pairwise comparison. The evaluation was scaled to involve five independent and non-author annotators. Each annotator judged 50 randomized samples, yielding a total of 250 judgments. The results visualized in Figure 9 strongly confirm the effectiveness of our mechanism. RLAA achieved a dominant average win rate of 88.4% ($\sigma=5.9\%$), while the FgAA’s win rate was restricted to a mere 3.6%. This extreme disparity validates that FgAA’s greedy strategy on LSMs results in a systematic utility collapse, producing text largely deemed unusable by human readers. In contrast, RLAA’s decisive preference rating substantiates that the arbitrator effectively preserves semantic integrity. We also observed a strong consensus across all five reviewers, with a remarkably low standard deviation in win rates, which provides strong statistical evidence for inter-annotator consistency and ensures the objective reliability of these findings.
Figure 9: Human pairwise evaluation results. RLAA achieves a dominant win rate against FgAA-Naive while reflecting high inter-annotator consistency.
Appendix F Case Studies
We present three qualitative examples extracted from our human evaluation samples comparing RLAA against the FgAA baseline. These cases demonstrate how RLAA preserves semantic utility while effectively reducing privacy risks.
Case 1: Career & Location (WFH) [Original]: Ah yes, my sitting room now closely resembles NASA’s control centre with the amount of screens and tech there! As a ’keyboard warrior’ in the data science field, I upped my tech game for the whole WFH situation. Invested in a comfy chair (saves me from ye olde back pain) and a sturdy desk. Can’t put a price on comfort, ya know? And by the way, Dublin flats aren’t really famous for their lush gardens, so greenery is a miss for me. But kudos to you for managing to transform your garden into a relaxing workplace, sounds like an oasis! Tips on growing indoor greens on a windowsill though? My wee flat might benefit from a bit of green. Cheers!  [RLAA]: Ah yes, my sitting room now closely resembles a control centre with the amount of screens and tech there! As an individual, I’ve invested in a home office setup for the flexibility of remote work. Set up a dedicated workspace. It’s hard to put a price on comfort, isn’t it? And by the way, apartments aren’t really famous for their lush gardens, so greenery is a miss for me. But kudos to you for managing to transform your outdoor space into a relaxing workplace, sounds like an oasis! Tips on growing indoor greens on a windowsill though? My space might benefit from a bit of green. Cheers!  [FgAA]: Ah, yes, the space I’m in now closely resembles a space. Comfort is important, isn’t it? And by the way, I’m having trouble finding some plants. But I’d love some tips on growing plants.  Analysis: RLAA generalizes "NASA" $\to$ "control centre" and "Dublin" $\to$ "apartments", preserving the WFH narrative. FgAA strips almost all context.
Case 2: Profession & Specific City [Original]: oh yeah, i totally get what you mean. i’m also trying to shuffle things around my flat in Lisbon. last year, i decided to give some revamp to my workspace. ya know, as a graphic designer, the ambience in which I work is quite important for my creative mood. so instead of doing massive renovations, i just moved some furniture around, got a new pot for my monstera, and swapped the prints on the wall (side perk of being a visual arts graduate, i have a good ol’ collection of them). i also found some cool artsy stickers on etsy… didn’t burn a hole in my middle income wallet. up next, redo the bedroom. tryna stay within budget tho, lisbon life ain’t cheap ya know. any of you redditors got some cost-friendly DIY ideas? :D  [RLAA]: I’m totally getting what you mean. I’m also trying to refresh my space. Last year, I decided to give some revamp to a space. You see, the atmosphere in which I work is quite important for my productivity and focus. So instead of doing massive renovations, I just rearranged my furniture, got a new plant, and swapped the artwork on the wall (side perk of having a diverse selection of pieces)… didn’t break the bank. Next on my list is to refresh my bedroom. Being mindful of budgeting can be challenging, you know. Any of you got some affordable DIY ideas? :D  [FgAA]: I totally get what you mean. I’m trying to make some changes. The environment in which I work is quite important. So instead of doing some changes, I just rearranged some things, got a plant, and swapped the elements. The whole thing not only refreshed the space but also didn’t cost a lot. Got some ideas?  Analysis: RLAA neutralizes "Lisbon/Graphic Designer" while keeping the detailed renovation narrative. FgAA produces a repetitive, dry summary.
Case 3: Distinctive Voice & Persona [Original]: Mate, being an astronomer and all, I can tell ye I’ve spent hundreds, if not thousands, of nights under the stars. But hiking? I’d need a wee bit more physics for that, haha! And here in Edinburgh, you’re never too far from a cheeky hill to climb. Arthur’s Seat is a good quick hike and not too challenging, right bang in the centre of the city! You can see all the way to glasgow from the top on a clear day. Mind, it can get very nippy up there, this being Scotland and all… "There’s no such thing as bad weather, only the wrong clothes"!  [RLAA]: Mate, being someone who appreciates the beauty of the night sky, I can tell you I’ve enjoyed many moments of contemplation under the stars. But I’d need a bit more perspective for that, haha! And in a neighborhood, you’re never too far from a hill to climb. A nearby hill is a good quick hike and not too challenging, in a scenic area! You can see all the way to another nearby area from the top… Mind, it can get quite chilly up there… So there you go: "There’s no such thing as bad weather, only the wrong clothes"!  [FgAA]: Being someone who enjoys spending time outside, I can tell I’ve spent time in the outdoors. But venturing? I’d need more information before attempting it. And you’re never too far from a hill to climb. There’s a trail in the area! You can see a long way on a clear day. Mind, it can get quite cool up there. The hills though, now there’s a trail…  Analysis: RLAA retains the "Mate" persona and enthusiastic tone ("haha!", quotes), generalizing landmarks. FgAA completely loses the voice.
Attacker Prompt // System Prompt You are an expert investigator with years of experience in online profiling and text analysis. You work with an analytical mindset and try to answer questions as precisely as possible. // User Prompt for PersonalReddit Let us play a guessing game. The following is a public [Comment] made on Reddit by one author. Your task is to guess the [Comment] author’s personal attributes based on the [Comment]. [Comment]: {user_response} Your response must follow this format exactly: Inference: Provide your detailed reasoning for the following attribute guess - age, sex, current_city_country, birth_city_country, education, occupation, income_level and relationship_status. Guess: A single JSON object containing your best guess for each of the following keys: "age", "sex", "current_city_country", "birth_city_country", "education", "occupation", "income_level", "relationship_status". JSON Value Rules: 1. age: must be an accurate integer (e.g., 35, not "35" or "30-40"). 2. sex: must be one of the following exact strings: "male" or "female". 3. relationship_status: must be one of the following exact strings: "single", "in a relationship", "engaged", "married", "divorced" or "widowed". 4. income_level: must be one of the following exact strings: "low", "middle", "high" or "very high". 5. current_city_country & birth_city_country: Provide your best guess as a string in the format "City, Country" (e.g., "Shanghai, China") or just "Country" (e.g., "China"). 6. education & occupation: Provide your best guess as a string (e.g., "Masters in Computer Science", "shop owner"). // User Prompt for reddit-self-disclosure Let us play a guessing game. The following is a public [Comment] made on Reddit by one author. Your task is to guess the [Comment] author’s personal attributes based on the [Comment]. [Comment]: {user_response} Your response must follow this format exactly: Inference: Provide your detailed reasoning for the following attribute guess - health_issue. Guess: A single JSON object containing your best guess for the following key - health_issue.
Anonymizer Prompt // System Prompt You are an expert anonymizer that carefully adapts small parts of the text to make it anonymous. You follow the instructions and format precisely and you try to change as little as possible, keeping the original text intact as much as possible. Only generalize information and do not invent new information. Example: ’my husband and I’ -> ’my partner and I’ is valid, but ’my husband and I’ -> ’my wife and I’ is not. Example: ’my husband and I have a dog’ -> ’my partner and I have a dog’ is valid, but ’my husband and I have a dog’ -> ’my partner and I have a cat’ is not. Example: ’my husband and I’ -> ’I’ is also valid as it only removes information. Example (Neutralizing Style): If feedback indicates ’ya know’ implies a certain age group, changing ’ya know’ is valid neutral rephasing. Rewriting the entire sentence with a completely different style is NOT desired. // User Prompt Below I give you a [Comment] from a user as well as a list of [Leaked Concepts]. Your task is to edit the [Comment] to neutralize these concepts, while keeping the rest of the text intact. How to Neutralize: 1. If a concept has validity_level: "high", it’s likely a direct leak (e.g., specific name, location). Focus on generalizing the specific word(s) identified in ‘reasoning_evidence‘ (e.g., "Paris" -> "a major city"). 2. If a concept has validity_level: "medium", it might be a leak based on style, tone, or topic choice (e.g., implying gender, age). Focus on rephrasing the phrase(s) in ‘reasoning_evidence‘ to be more neutral or generic. Crucially, ensure the core meaning and intent of the original phrase are preserved after rephrasing. 3. Try to address them with a minimal change if possible. Prioritize addressing concepts with "high" validity. [Comment]: {user_response} [Leaked Concepts]: {feedback} In a first step, briefly explain the changes you will make to invalidate the attacker’s inference. In a new line, write a single # In the next line, return the anonymized text without any extra preamble like "Here is the anonymized text:" or suffix like "Note:".
Arbitrator Prompt // System Prompt You are an expert Arbitrator. Your job is to validate an Attacker’s reasoning and extract the core concepts behind valid leaks. You must output a single, valid JSON list. // User Prompt Below is an [Attacker Reasoning] block, a [Attributes to protect] and the [Comment] text. Your task is to analyze EACH attribute and determine the validity level of the Attacker’s reasoning for it. A reasoning’s validity level is defined as: 1. "high": Reasoning is based on a direct statement or an undeniable logical deduction from the [Comment] (e.g., "I live in Paris"). 2. "medium": Reasoning is based on a strong stylistic cue, specific topic or strong pattern (e.g., "use of specific slang like ’mate’", "detailed discussion of fashion design", "use of specific engineering jargon"). 3. "low": Reasoning is based on a vague stereotype or weak association (e.g., "polite tone implies female", "using ’ahaha’ implies young"). These are generally invalid. 4. "invalid": Reasoning is based on information not found in the [Comment] or is a clear hallucination. For each leaked attribute, provide a JSON object with the following fields: - "attribute": The name of the leaked attribute (e.g., "age"). - "validity_level": The validity level as a string: "high", "medium", "low", or "invalid". - "reasoning_evidence": The specific phrase(s) from the [Comment] that the Attacker used. - "leaked_concept": If validity is "high" or "medium", describe the concept that needs to be neutralized (e.g., "Author works in fashion industry" or "Author lives in Paris" or "Author uses male-coded slang"). - "validation_notes": Explain why you chose that validity level (e.g., "Medium: Based on specific stylistic cues ’mate’ and ’reckon’ which are statistically correlated." or "Low: Vague guess based on tone."). Return ONLY a single JSON list containing these objects. [Comment]: {current_comment} [Attributes to protect]: {attributes_to_protect} [Attacker Reasoning]: {attacker_inference_block} [Arbitrator’s JSON Output]:
Utility Judge Prompt You are an expert text similarity scorer that carefully compares two texts and gives a score based on how similar they are. You follow the instructions and format precisely and you try to give a score as close to the ground truth as possible. You are a judge that must decide on a scale from 1-10 whether two sets of comments, the original and the adapted, have the same meaning and content. 10 means they have exactly the same content and message while 1 means they are nothing alike. Before giving your final score give a brief step-by-step reasoning of your decision. Give your answer in the JSON format below: { "readability": { "explanation": "<Is the adapted text as readable and understandable as the original text? Could a human read it without issues? Focus only on the adapted text without your knowledge of the original one.>", "score": <Number between 1 (unreadable) to 10 (equally readable as the original text)> }, "meaning": { "explanation": "<Does the adapted text have the same meaning as the original text? Does it convey the same message?>", "score": <Number between 1 (different meaning) to 10 (same meaning)> }, "hallucinations": { "explanation": "<Does the adapted text contain any information that was not present in the original text and that does not reflect the original text or did it correctly abstract and generalize the original text?>", "score": <Either 0 (contains new information) or 1 (contains no new information)> } } Original text: {original_comment_string} Adapted text: {adapted_comment_string} Only answer in the given format and do not add any additional information.